### v1

# A free boundary problem in accretive growth

Abstract.
We prove an existence result for a free boundary problem inspired by the modelization of accretive growth. The growth process is formulated through a level-set approach, leading to a boundary-value problem for a Hamilton–Jacobi equation within a prescribed constraining set. Existence, variational representability, and regularity of solutions to the growth subproblem are investigated. The full system arises from coupling the growth dynamics with an elliptic equation for the activation field. Existence of solutions to the fully coupled free boundary problems is obtained via an iterative procedure.
Key words and phrases:
Accretive growth, Hamilton–Jacobi equation, viscosity solution, John domain, free boundary problem.
2020 Mathematics Subject Classification:
35R35, 35F21, 35F30, 35D40
1. Introduction
We are interested in the free boundary problem
$H(x,Ku(x,v(x)),\nabla v)=0\quad\text{in}\ \Omega\setminus\overline{V_{0}},$ (1.1)
$-\Delta u(x,t)=1\quad\text{for}\ x\in\{v<t\},\ t>0.$ (1.2)
Here, $\Omega\subset{\mathbb{R}}^{n}$ is an open set, possibly being unbounded, and $V_{0}\subset\Omega$. The unknowns are the two scalar fields $v:\Omega\to[0,\infty)$ and $u:Q_{v}\to{\mathbb{R}}$, where the latter is defined on the a-priori unknown space-time noncylindrical domain $Q_{v}:=\{(x,t)\in{\mathbb{R}}^{n}\times(0,\infty)\ |\ v(x)<t\}$. The Hamiltonian $H:\overline{\Omega}\times{\mathbb{R}}\times{\mathbb{R}}^{n}\to{\mathbb{R}}$ is continuous, with $p\mapsto H(x,u,p)$ having compact, convex, and nondegenerate sublevels, and $K$ is a space-time convolution operator.
The main result of the paper, Theorem 2.1, states that there exists a solution $(v,u)$ to (1.2)–(1.1) under mixed Dirichlet-Neumann conditions. The stationary Hamilton–Jacobi equation (1.1) is solved in the viscosity sense, while the elliptic equation (1.2) is treated in the weak sense. Along the way, we also discuss existence, variational representability, and regularity of sublevels for viscosity solutions of (1.1) for a given $u$, see Proposition 3.3 and Proposition 3.8.
1.1. Accretive growth
Problem (1.1)–(1.2) is inspired by the modelization of an accretive-growth process in a confined environment. Accretive growth occurs across a variety of different biological systems, as well as in many natural and technological settings. The formation of horns, teeth and seashells [48, 54], coral reefs [35], bacterial colonies [30], trees [25], and cell motility due to actin growth [29] are biological examples of accreting systems, among many others [28]. In geophysics, sedimentation [12] and glacier evolution [32] are also accretive processes, so as planet formation [45]. Furthermore, accretive growth is a key aspect in metal solidification [46], crystal growth [38], additive manufacturing [31], layering, coating, and masonry, just to mention a few.
The growth problem is assumed to be constrained to a given smooth set $\overline{\Omega}\subset{\mathbb{R}}^{n}$, which may be chosen to be unbounded. We follow the classical level-set approach [24, 52]. Starting from the initial shape $V_{0}\subset\Omega$, the growing system is described by the set-valued function $t\mapsto V(t)\subset{\mathbb{R}}^{n}$, whose values are the sublevels of the unknown function $v:\overline{\Omega}\to[0,\infty)$, namely,
$V(t):=\{x\in\overline{\Omega}\ |\ v(x)<t\}\quad\text{for}\ t>0.$
The value $v(x)$ models the time at which the point $x\in\overline{\Omega}$ is added to the growing system. Correspondingly, $v$ is called time-of-attachment function in the following.
Growth is always influenced by its environment and aspects such as nutrient or oxygen concentration, pressure, temperature, or stress, [28] have often to be taken into account. We model this fact by considering the simplest situation of a scalar activation field $u:Q_{v}\to{\mathbb{R}}$ driving the growth. The dynamics of the activation field $u$ is influenced by the actual shape of the growing system. Since growth is often very slow with respect to other processes, we assume that $u$ evolves quasistatically. We model a basic quasistatic diffusion process by the linear elliptic equation (1.2) on the growing domain $V(t)$ for all $t>0$. This corresponds to the case of $u$ being a concentration or temperature in an homogeneous medium and could of course be generalized. In particular, the choice of the right-hand side equal to $1$ is made solely for simplicity. Note that the resulting full model (1.1)–(1.2) is of free boundary type.
Starting from the initial shape $V_{0}$, the growing system $V(t)$ may be assumed to evolve by layering material at its boundary at a velocity depending on the position $x\in\partial V(t)$, on the outward-pointing normal direction $n(x)$ to $\partial V(t)$, and on the activation field $u$. The evolution in time of a point $x(t)\in\partial V(t)$ is modeled by the ODE
$x^{\prime}(t)=f\big(x(t),u(x(t),t),n(x(t))\big),$
where $f:{\mathbb{R}}^{n}\times{\mathbb{R}}\times{\mathbb{R}}^{n}\to{\mathbb{R}}^{n}$ is a given velocity field and the prime indicates the time derivative. Assuming smoothness of $\partial V(t)$ (an assumption which we will soon be forced to abandon, however), differentiating $t=v(x(t))$ with respect to time, and setting $n(x)=\nabla v(x)/|\nabla v(x)|$, the function $v$ is then asked to satisfy $H(x,u(x,v(x)),\nabla v(x))=0$ for the Hamiltonian $H(x,u,p)=p\cdot f(x,u,p/|p|)-1$. A reference situation is that of normal accretion, where $f(x,u,n)=\gamma(x,u,n)n$ for some scalar accretion rate $\gamma:{\mathbb{R}}^{n}\times{\mathbb{R}}\times{\mathbb{R}}^{n}\to(0,\infty)$. In this case, the Hamiltonian reads $H(x,u,p)=\gamma(x,u,p/|p|)|p|-1$. We make no assumption on the specific form of the Hamiltonian in the following, but simply ask $H$ to be continuous and the sets $\{p\in{\mathbb{R}}^{n}\,|\,H(x,u,p)\leq 0\}$ to be convex, bounded, and nondegenerate.
A crucial trait of the model is the occurrence of the term $u(x,v(x))$ into the Hamiltonian $H$ in (1.1). Indeed, this reflects the assumption that the growth process is driven by the value of the activating field at the growing front. On the other hand, treating the composition $x\mapsto u(x,v(x))$ of the two unknowns $u$ and $v$ into the Hamiltonian $H$ poses severe mathematical challenges. In fact, $V(t)$ may develop singularities, see Figure 1, and can change topology. As a consequence, the activation field $u$ cannot be assumed to be regular in time, which prevents from taking its trace on the a-priori unknown and nonsmooth space-time manifold $\{(x,t)\in{\mathbb{R}}^{n}\times(0,\infty)\ |\ t=v(x)\}$. We are hence forced to regularize the problem by replacing $u$ in the Hamilton–Jacobi equation (1.1) by a space-time convolution $Ku$. This provides the additional regularity needed to take the trace of the field $u$ on $\{(x,t)\in{\mathbb{R}}^{n}\times(0,\infty)\ |\ t=v(x)\}$. From the modeling viewpoint, this relaxation implies that the growth is driven by a local mean of the values of the activating field $u$ at the accreting front, instead of its pointwise ones. This modification has a limited impact on the modeling, as we are entitled to choose the support of the regularizing kernels to be arbitrarily small, so to localize the regularization.
1.2. Boundary conditions
Figure 1. Setting and notation.
The PDE system (1.1)–(1.2) calls for specifying boundary conditions. As $V_{0}\subset\Omega$ models the initial shape and $v$ is the time-of-attachment function, the Hamilton–Jacobi equation (1.1) is naturally complemented by the condition
$v=0\quad\text{on}\ V_{0}.$ (1.3)
At the boundary $\partial\Omega$ of the constraining set we can consider the boundary condition
$H(x,Ku(x,v(x)),\nabla v)\geq 0\quad\text{for}\ x\in\partial\Omega\setminus\overline{V_{0}}.$ (1.4)
In the setting of viscosity solutions to Hamilton–Jacobi constrained problems, this boundary condition is classical (and qualifies $v$ as constrained viscosity solution), see [14, 50, 51].
The activation field $u$ is subject to mixed homogeneous Dirichlet–Neumann conditions. To pose boundary conditions on $u$ is subtle, as $\partial V(t)$ need not be smooth. The Dirichlet condition is hence solely imposed on $\Gamma\subset\partial V_{0}\cap\partial\Omega$, which is assumed to be nonempty, open (with respect to the topology of $\partial\Omega$), and Lipschitz. As $V(t)$ grows with time, we have that $\Gamma\subset\partial V(t)$ for all $t>0$ and we can prescribe
$u=0\quad\text{on}\ \Gamma\times(0,\infty).$ (1.5)
The homogeneous Neumann condition is then imposed on the complement $\partial V(t)\setminus\overline{\Gamma}$ (the closure is taken with respect to the topology of $\partial\Omega$), namely,
$\partial u/\partial n=0\quad\text{on}\ \partial V(t)\setminus\overline{\Gamma}\times(0,\infty).$ (1.6)
Note nonetheless that condition (1.6) will be imposed just weakly: We address a variational formulation of equation (1.2), see (2.12) below, ensuring the validity of the Neumann boundary condition where $\partial V(t)$ is smooth enough.
Before moving on, let us remark that the Neumann boundary condition (1.6) is also imposed on the portion $A(t):=(\partial V(t)\setminus\overline{\Gamma})\cap\partial\Omega$, which is nonempty, see Figure 1. From the modeling viewpoint, this is somehow still unsatisfactory, as one would prefer to have the Dirichlet condition there. This would however require $A(t)$ to be sufficiently regular in order for a trace to be well defined and such regularity of $A(t)$ is presently not known.
1.3. Results
Our main result, Theorem 2.1, establishes the existence of suitably weak solutions $(v,u)$ to (1.1)–(1.6). More precisely, the Hamilton–Jacobi equation (1.1), subject to conditions (1.3)–(1.4), is solved in the viscosity sense, whereas the elliptic equation (1.2) on the unknown domain $V(t)$, under conditions (1.5)–(1.6), is treated variationally.
Existence of solutions is proved by an iterative argument, where we alternately solve the Hamilton–Jacobi subproblem (1.1), (1.3)–(1.4) for a given $u$, and the elliptic subproblem (1.2), (1.5)–(1.6) for a given $v$.
In addition to accretive growth, the Hamilton–Jacobi subproblem (1.1), (1.3)–(1.4) (for $u$ given) arises in different contexts, from front propagation and geometric optics [6, 24, 52], to optimal control [4], differential games [15, 27], and image reconstruction [3]. As a consequence, it has attracted early attention, see for instance the classical references [14, 19, 40]. Different results are available, corresponding to different assumptions on the ingredients of the problem. Still, we have not been able to find result which exactly cover the specific framework of (3.5)–(3.9) below. For this reason, we present complete arguments, without necessarily claiming originality. Existence and variational representation of nonnegative viscosity solutions to (1.1), (1.3)–(1.4) (for $u$ given) are proved in Proposition 3.3.
The solution to the elliptic subproblem (1.2), (1.5)–(1.6) for a given $v$ relies on the possibility of establishing a Poincaré inequality on $V(t)$. In turn, this requires proving that $V(t)$ maintains some minimal level of regularity. In Proposition 3.8, we show that, starting from some suitably regular $V_{0}$, the growing set $V(t)$ is a John domain for all times (see Section 2). This extends to the constrained case $\Omega\not={\mathbb{R}}^{n}$ the arguments in [22]. Some technical modifications are required for this extension, since the boundary $\partial\Omega$ of the constraining domain must also be taken into account. Note that John regularity is sharp is this setting [22].
Eventually, the convergence of the iterative scheme depends on the time regularity of the set-valued maps $t\mapsto V(t)$ for different $v$. These happen to be Lipschitz continuous with respect to the Hausdorff distance, as well as Hölder continuous with respect to Lebesgue measure of the differences. Crucially, such continuities are uniform with respect to the iteration, allowing for the use of compactness arguments.
Let us conclude this section by noting that uniqueness of solutions to the coupled system (1.1)–(1.6) remains open. Even under the boundary condition (1.4), classical uniqueness arguments [5, 14, 40] seem to be not applicable. In fact, without further assumptions uniqueness seems to fails even in the simplest decoupled setting of $H$ not depending on $u$, see Remark 3.5 below.
Before moving on, let us mention that the results of this paper can be easily generalized to more complex settings. A first generalization consists in considering the Hamilton–Jacobi equation
$H(x,v(x),u(x,v(x)),\nabla v)=0\quad\in\text{in}\ \Omega\setminus\overline{V_{0}}$
where an additional dependence on $v$ is included. This extension would allow to cover the case of a time-dependent growth process, though at the cost of a somewhat more intricate analysis. A second possible direction is to replace equation (1.2) with other stationary PDEs, provided the driving operator retains suitable compactness and continuity properties. In particular, (1.1) can be replaced by a semilinear elliptic system with sufficiently controlled nonlinearities.
1.4. Connections with previous studies
The theory of the Hamilton–Jacobi subproblem (1.1) with condition (1.3) for $u$ given (and suitably smooth) is well developed. We limit ourselves in mentioning the classical references [4, 5, 19, 40] for the general frame of viscosity solutions. In the context of optimal control with state constraints, condition (1.4) corresponds to the notion of constrained viscosity solution, going back to [14, 50, 51], see also [1, 26, 33, 36, 41] among many others. A class of biological invasion models based on Hamilton–Jacobi dynamics with obstacles has been introduced in [9] and studied in [10, 23], see also [7, 11] and the reference therein. Our results rely in a fundamental way on the metric structure underlying the Hamilton–Jacobi subproblem, which is already pointed out in [18, 40], see also [44, 47]. A wide range of results on the Hamilton–Jacobi subproblem is already available. However, to the best of our knowledge, none of them directly addresses the specific assumptions adopted here, see (3.5)–(3.9). For this reason, we provide detailed proofs.
When considering the full free boundary problem, the closest paper to the present one is [21]. This focuses on (1.1)–(1.2) for the specific generalized eikonal case $H(x,u,p)=\gamma(u)|p|-1$ and for (1.2) replaced by the linearized elastic equilibrium system. The setting in [21] is however much simplified by the absence of a constraining set, i.e., $\Omega={\mathbb{R}}^{n}$, and by the choice the docking condition $u=0$ on $\omega\times(0,\infty)$ with $\omega\subset V_{0}$ instead of the Dirichlet condition (1.5). The same docking condition could be considered here, as well, and would lead to a simpler theory. Note nonetheless that the applicative relevance of such a docking condition is limited to $n=2$, as it cannot be practically realized for $n\geq 3$.
A related set of questions on tumor-growth modeling and Hamilton–Jacobi dynamics is addressed by Collins, Jacobs, & Kim in [17], although in the context a different model. There, the focus lies on the regularity of the free boundary of a tumor model driven by pressure and by nutrient concentration. The main character of the analysis is again the function $v$, which is proved to be Hölder continuous with regular sublevels, except possibly on a subset of Hausdorff dimension $n-\mu$ for some $\mu\in(0,1)$. Although the underlying dynamics essentially differ, these results bear a close analogy to Proposition 3.8 below. A previous analysis of the same model in specific parameter regimes is in [34, 37].
More generally, one can refer to the extensive literature on Hele–Shaw/Muskat/Stefan dynamics, see, e.g., [13, 55]. In these models, however, growth is driven by a fundamentally different underlying mechanism. In the Hele–Shaw setting, the driving field satisfies a Dirichlet boundary condition, and the domain evolves with a normal velocity proportional to the gradient of this field at the free boundary. In contrast, in our framework the activating field $u$ u is subject to a homogeneous Neumann condition, while the velocity of the free boundary depends on the Dirichlet values of $u$ at the growing front.
[…]
1.5. Structure of the paper
In Section 2, we introduce notation, specify assumptions, and state the main result, i.e., Theorem 2.1. Section 3 is devoted to the study of the Hamilton–Jacobi subproblem. Eventually, the proof of Theorem 2.1 is presented in Section 4.
2. Setting and main result
In this section, we fix the assumptions on the ingredients of the problem, present the notion of solution, and state our main result, Theorem 2.1.
In the following, we use ${\mathbb{R}}_{+}=(0,\infty)$, denote by $B_{r}(x)$ the open ball with center $x\in{\mathbb{R}}^{n}$ and radius $r>0$, and simply write $B_{r}$ for $B_{r}(0)$. We use the same symbol $|\cdot|$ for the Euclidean norm in ${\mathbb{R}}^{m}$ for $m\geq 1$ and for the Lebesgue measure of any measurable set. We let $\overline{A}$, $A^{\circ}$, $A^{c}$, and $1_{A}$ be the closure, the interior, the complement in ${\mathbb{R}}^{n}$, and the characteristic function of a set $A\subset{\mathbb{R}}^{n}$, respectively. Specifically, $1_{A}(x)=1$ if $x\in A$ and $1_{A}(x)=0$ if $x\in A^{c}$. We indicate by $d(x,A)=\inf_{a\in A}|x-a|$ the distance to a nonempty set $A\subset{\mathbb{R}}^{n}$. For any two sets $A,\,B\subset{\mathbb{R}}^{n}$, the symbol $A+B$ stands for the usual sum. In case $A$ and $B$ are nonempty, $d_{H}(A,B)$ denotes their Hausdorff distance, namely $d_{H}(A,B):=\max\{e(A,B),e(B,A)\}$ where $e(A,B):=\sup_{a\in A}d(a,B)$ is the excess of $A$ with respect to $B$.
2.1. Assumptions
The growth process is constrained to a given
$\Omega\subset{\mathbb{R}}^{n}\quad\text{nonempty, open, connected, and $C^{1,1}$}.$ (2.1)
This regularity implies that all points $x\in\partial\Omega$ can be touched by an internal and by an external ball with radius $\overline{\varepsilon}>0$, namely,
$\exists\overline{\varepsilon}>0\ \forall x\in\partial\Omega\ \exists x_{i},\,\exists x_{e}\in{\mathbb{R}}^{n}:$
$\quad B_{\overline{\varepsilon}}(x_{i})\subset\Omega,\quad B_{\overline{\varepsilon}}(x_{e})\subset({\overline{\Omega}})^{c},\quad\text{and}\ \ |x-x_{i}|=|x-x_{e}|=\overline{\varepsilon},$
see [39, Cor. 2]. This entails that the projection $\pi:\partial\Omega+B_{\overline{\varepsilon}}\to\partial\Omega$ given by
$\pi(x):={\rm arg\,\min}_{y\in\partial\Omega}|x-y|,$
as well as the normal outward-pointing vector field $\nu:\partial\Omega\to{\mathbb{S}}^{n-1}$ are well-defined and Lipschitz continuous. In the following, the Lipschitz continuous composition $\eta:=-\nu\circ\pi:\partial\Omega+B_{\overline{\varepsilon}}\cap\overline{\Omega}\to{\mathbb{S}}^{n-1}$ will play a role in the proof of Proposition 3.8.
As $\Omega$ is path-connected and $C^{1,1}$, its closure $\overline{\Omega}$ is path-connected, as well. In particular, $\overline{\Omega}$ is a geodesic space. For later use, we indicate by $d_{\overline{\Omega}}$ the geodesic distance on $\overline{\Omega}$, namely,
$d_{\overline{\Omega}}(x,y)=\min\left\{\int_{0}^{1}|\gamma^{\prime}(s)|\,{\rm d}s\ \Big|\ \gamma\in W^{1,\infty}(0,1;\overline{\Omega}),\ \gamma(0)=y,\ \gamma(1)=x\right\}\quad\forall x,\,y\in\overline{\Omega}.$
Given $x,\,y\in\overline{\Omega}$, we call geodesic from $y$ to $x$ any curve realizing the minimum above, recalling Note that geodesics are not necessarily unique. We additionally ask that the Euclidean and the geodesic distance are comparable in $\overline{\Omega}$, namely,
$\exists L>0:\quad d_{\overline{\Omega}}(x,y)\leq L|x-y|\quad\forall x,\,y\in\overline{\Omega}.$ (2.2)
Note that this does not follow from $C^{1,1}$ regularity, as $\Omega$ may be unbounded.
Let us recall that a bounded set $U\subset{\mathbb{R}}^{n}$ is said to be a John domain (or John regular) if there exists a point $x_{0}\in U$ (John center) and a constant $\kappa\in(0,1]$ (John constant) such that for all points $x\in U$ one can find an arc-length parametrized curve $\gamma:[0,\ell_{\gamma}]\to U$ with $\gamma(0)=x$, $\gamma(\ell_{\gamma})=x_{0}$, and $d(\gamma(s),\partial U)\geq\kappa s$ for all $s\in[0,\ell_{\gamma}]$, see [42]. In other words, the twisted cone $\cup_{s\in[0,\ell_{\gamma}]}B_{\kappa s}(\gamma(s))$ is asked to be contained in $U$. Note that John domains are connected.
On the initial set $V_{0}$ we ask the following
$V_{0}\subset\Omega\quad\text{is nonempty, open, and John regular with respect}$
$\quad\text{to $x_{0}\in V_{0}$ with John constant $\kappa_{0}\in(0,1]$}.$ (2.3)
Note that, given (2.3), $\overline{V_{0}}$ is a John domain with respect to $x_{0}$ with John constant $\kappa_{0}$, as well. In the following, we indicate by ${\rm Lip}_{x}$ the set of Lipschitz curves connecting $\overline{V_{0}}$ to $x\in{\mathbb{R}}^{n}$, namely,
${\rm Lip}_{x}=\{\gamma\in W^{1,\infty}(0,1;{\mathbb{R}}^{n})\ |\ \gamma(0)\in\overline{V_{0}},\ \gamma(1)=x\}\quad\forall x\in{\mathbb{R}}^{n}.$
In order to specify boundary conditions for (1.2) we assume the following
$\Gamma\subset\partial V_{0}\cap\partial\Omega\ \text{nonempty, open in $\partial\Omega$, with $d_{H}(\Gamma,\partial V_{0}\setminus\partial\Omega)>0$}.$ (2.4)
As $V(t)\subset\Omega$ grows with $t$, assumption (2.4) guarantees that $\Gamma\subset\partial V(t)\cap\partial\Omega$ for all $t>0$ and $d_{H}(\Gamma,\partial V(t)\setminus\partial\Omega)$ is nondecreasing with respect to $t$.
As already mentioned, the open set $V(t)$ could be not smooth, see Proposition 3.8. Nonetheless, under assumptions (2.1) and (2.4), the portion $\Gamma$ of $\partial V(t)$ is smooth and is well separated from $\partial V(t)\setminus\partial\Omega$ for all $t>0$. This allows to define the trace on $\Gamma$ of any $u\in H^{1}(V(t))$. More precisely, set $G\subset\Gamma+B_{\delta}$ for some $\delta<d_{H}(\Gamma,\partial V_{0}\setminus\partial\Omega)$ small enough with $G\subset\subset V_{0}$ Lipschitz. Then, the trace operator on $H^{1}(V(t))$ can be defined by composing the restriction $u\in H^{1}(V(t))\mapsto u|_{G}\in H^{1}(G)$ and the trace operator $\gamma:H^{1}(G)\to H^{1/2}(\Gamma)$. This allows to specify
$H^{1}_{\Gamma}(V(t)):=\{u\in H^{1}(V(t))\ |\ u=0\ \text{on}\ \Gamma\}\quad\text{for}\ t>0.$
The Hamiltonian $H$ is asked to fulfill the following
$H:\overline{\Omega}\times{\mathbb{R}}\times{\mathbb{R}}^{n}\to{\mathbb{R}}\quad\text{is continuous},$ (2.5)
$C_{xu}:=\{p\in{\mathbb{R}}^{n}\ |\ H(x,u,p)\leq 0\}\quad\text{is convex}\quad\forall(x,u)\in\overline{\Omega}\times{\mathbb{R}},$ (2.6)
$C_{xu}=\overline{\{p\in{\mathbb{R}}^{n}\ |\ H(x,u,p)<0\}}\quad\forall(x,u)\in\overline{\Omega}\times{\mathbb{R}},$ (2.7)
$\forall(x,u,p)\in\overline{\Omega}\times{\mathbb{R}}\times{\mathbb{R}}^{n}:\ \exists\lambda\in(0,1)\ \text{such that}\ p/\lambda\in C_{xu}\ \Rightarrow\ H(x,u,p)<0,$ (2.8)
$\exists 0<\sigma_{*}\leq\sigma^{*}:\quad B_{\sigma_{*}}\subset C_{xu}\subset B_{\sigma^{*}}\quad\forall(x,u)\in\overline{\Omega}\times{\mathbb{R}}.$ (2.9)
Note that $p\mapsto H(x,u,p)$ is not asked to be convex but merely to have convex sublevels with respect to $p$, see (2.6). Assumptions (2.7)–(2.8) are of topological nature and control the size of $\partial C_{xu}$. In particular, given the continuity assumption in (2.5), property (2.8) follows if $H(x,u,0)<0$, without being equivalent (take H(x,u,p)=(|p|1)d(p,A)H(x,u,p)=(|p|-1)d(p,A) for $A=\partial B_{1}\cup\{t(\cos 2\pi t,\sin 2\pi t)\>|\>t\in[0,1]\}\in{\mathbb{R}}^{2}$). The simplest setting where (2.5)–(2.9) hold is that of the generalized eikonal equation, namely, $H(x,u,p)=\gamma(x,u)|p|-1$ with $\gamma$ continuous and $1/\sigma^{*}\leq\gamma(\cdot)\leq 1/\sigma_{*}$.
For all $(x,u)\in\overline{\Omega}\times{\mathbb{R}}$ we define the support function $q\in{\mathbb{R}}^{n}\mapsto\sigma(x,u,q)$ of the nonempty, convex, and closed set $C_{xu}$ as
$\sigma(x,u,q)=\sup\{q\cdot p\ |\ H(x,u,p)\leq 0\}.$
The functions $q\in{\mathbb{R}}^{n}\mapsto\sigma(x,u,q)$ are nonnegative and positively $1$-homogeneous. In particular, $0=\sigma(x,u,0)=\min_{q}\sigma(x,u,q)$.
The relaxation kernels are asked to fulfill
$k\in W^{1,1}({\mathbb{R}}_{+}),\quad\phi\in H^{1}({\mathbb{R}}^{n}).$ (2.10)
Given $v\in C(\overline{\Omega})$, we set $V(t)=\{x\in\overline{\Omega}\ |\ v(x)<t\}$ and $Q_{v}:=\cup_{t>0}V(t)\times\{t\}$. Let us note that $Q_{v}=\{(x,t)\in\overline{\Omega}\times{\mathbb{R}}_{+}\ |\ v(x)-t\leq 0\}$ with $(x,t)\in\overline{\Omega}\times{\mathbb{R}}_{+}\mapsto v(x)-t$ continuous. In particular, $Q_{v}\subset\overline{\Omega}\times{\mathbb{R}}_{+}$ is measurable.
For all $u:Q_{v}\to{\mathbb{R}}$ measurable with $u(\cdot,t)\in H^{1}_{\Gamma}(V(t))$ for a.e. $t>0$ and $t\mapsto\|u(\cdot,t)\|_{H^{1}_{\Gamma}(V(t))}\in L^{\infty}({\mathbb{R}}_{+})$ we define the space-time convolution
$Ku(x,t):=\int_{0}^{t}\!\!\int_{{\mathbb{R}}^{n}}k(t-s)\phi(x-y)\,\overline{u}(y,s)\,{\rm d}y\,{\rm d}s$
where $\overline{u}(\cdot,t)$ denotes the trivial extension of $u(\cdot,t)$ to the whole ${\mathbb{R}}^{n}$ for a.e. $t>0$. As $\overline{u}\in L^{\infty}({\mathbb{R}}_{+};L^{2}({\mathbb{R}}^{n};{\mathbb{R}}^{n}))$, by Young’s convolution inequality we get that $Ku$ is Lipschitz continuous in space-time. More precisely, from
$\nabla(Ku)_{t}(x,t)=k(0)\int_{{\mathbb{R}}^{n}}\nabla\phi(x-y)\,\overline{u}(y,t)\,{\rm d}y+\int_{0}^{t}\!\!\int_{{\mathbb{R}}^{n}}k^{\prime}(t-s)\nabla\phi(x-y)\,\overline{u}(y,s)\,{\rm d}y\,{\rm d}s,$
we readily check that
$\|Ku\|_{W^{1,\infty}({\mathbb{R}}_{+};W^{1,\infty}({\mathbb{R}}^{n};{\mathbb{R}}^{n}))}\leq c\sup_{t>0}\|\overline{u}\|_{L^{\infty}({\mathbb{R}}_{+};L^{2}({\mathbb{R}}^{n},{\mathbb{R}}^{n}))}$ (2.11)
where the constant $c>0$ depends on $\|k\|_{W^{1,1}({\mathbb{R}}_{+})}$ and $\|\phi\|_{H^{1}({\mathbb{R}}^{n})}$ only.
We are now ready to state our main result.
Theorem 2.1 (Existence for the free boundary problem).
Under assumptions (2.1)–(2.10) there exist $v\in C(\overline{\Omega})$ and $u:Q_{v}\to{\mathbb{R}}$ measurable with $u(\cdot,t)\in H^{1}_{\Gamma}(V(t))$ for a.e. $t>0$ such that
$\int_{V(t)}\nabla u\cdot\nabla w\,{\rm d}x=\int_{V(t)}w\,{\rm d}x\quad\forall w\in H^{1}_{\Gamma}(\Omega),\ \text{for a.e.}\ t>0,$ (2.12)
$v(x)=\min\Bigg\{\int_{0}^{1}\sigma\big(\gamma(s),Ku\big(\gamma(s),v(\gamma(s))\big),\gamma^{\prime}(s)\big)\ {{\rm d}s}\ \Big|\ \gamma\in{\rm Lip}_{x},\ \gamma(\cdot)\in\overline{\Omega}\Bigg\}\quad\forall x\in\overline{\Omega},$ (2.13)
and $v$ is a viscosity solution to (1.1), namely, for any $\varphi\in C^{1}({\mathbb{R}}^{n})$, we have that
$H(x,Ku(x,v(x)),\nabla\varphi(x))\leq(\geq)\,0$
at any local minimum (maximum, respectively) point $x\in\Omega\setminus\overline{V_{0}}$ of $\varphi-v$.
Moreover, by additionally asking
$\forall\delta>0\ \exists\lambda\in(0,1)\ \forall(x,u,p)\in\overline{\Omega}\times{\mathbb{R}}\times{\mathbb{R}}^{n}:\ H(x,u,p)<-\delta\Rightarrow p/\lambda\in C_{xu}$ (2.14)
instead of (2.7), the function $v$ is a constrained viscosity solution, namely, for any $\varphi\in C^{1}({\mathbb{R}}^{n})$, we have that $H(x,Ku(x,v(x)),\nabla\varphi(x))\geq 0$ at any local maximum point $x\in\partial\Omega\setminus\overline{V_{0}}$ of $\varphi-v$.
We remark that the minimum in (2.13) is always attained. Note that the function $v$ appears in the argument of the integrand in (2.13), as well. As such, (2.13) does not directly define $v$, given $u$. In fact, no uniqueness is implied by (2.13) in general. Moreover, note that the test function $w$ in (2.12) is taken in the whole $H^{1}_{\Gamma}(\Omega)$ instead of $H^{1}_{\Gamma}(V(t))$ for a.a. $t>0$. Given $w\in H^{1}_{\Gamma}(\Omega)$, its restriction $w|_{V(t)}$ belongs to $H^{1}_{\Gamma}(V(t))$ for all $t>0$. On the other hand, $V(t)$ need not be an extension domain, see Figure 1, so that a $\overline{w}\in H^{1}_{\Gamma}(V(t))$ may have no extension to $H^{1}_{\Gamma}(\Omega)$. Hence, satisfying (2.12) for $w\in H^{1}_{\Gamma}(\Omega)$ delivers a weaker statement with respect to asking it for $w\in H^{1}_{\Gamma}(V(t))$ for a.a. $t>0$. However, (2.12) still implies that $u$ solves the elliptic problem in distributional sense, and that the homogeneous Neumann condition at $\partial V(t)\setminus\Gamma$ is locally recovered, wherever the boundary $\partial V(t)$ is smooth enough.
Theorem 2.1 is proved in Section 4 via an iteration procedure.
3. Some remarks on the Hamilton–Jacobi subproblem
In preparation to the proof of Theorem 2.1, we collect some material on the Hamilton–Jacobi subproblem. More precisely, given $\Omega\subset{\mathbb{R}}^{n}$ fulfilling (2.1) and
$V_{0}\subset\Omega\quad\text{nonempty and open},$ (3.1)
in this section, we are interested in the solvability in the viscosity sense of problem
$H(x,\nabla v)=0\quad\text{in}\ \Omega\setminus\overline{V_{0}},$ (3.2)
$v=0\quad\text{in}\ V_{0}.$ (3.3)
We recall that $v\in C(\overline{\Omega})$ is said to be a viscosity solution to equation (3.2) if, for any $\varphi\in C^{1}({\mathbb{R}}^{n})$, we have that $H(x,\nabla\varphi(x))\leq(\geq)\,0$ at any local minimum (maximum, respectively) point $x\in\Omega\setminus\overline{V_{0}}$ of $\varphi-v$. A viscosity solution $v\in C(\overline{\Omega})$ to (3.2) is said to be a constrained viscosity solution if
$H(x,\nabla v)\geq 0\quad\text{on}\ \partial\Omega\setminus\overline{V_{0}}$ (3.4)
in the viscosity sense, namely, for any $\varphi\in C^{1}({\mathbb{R}}^{n})$, we have that $H(x,\nabla\varphi(x))\geq 0$ at any local maximum point $x\in\partial\Omega\setminus\overline{V_{0}}$ of $\varphi-v$.
In case $\Omega={\mathbb{R}}^{n}$, we refer to problem (3.2)–(3.3) as unconstrained, whereas, if $\Omega\not={\mathbb{R}}^{n}$ we say that problem (3.2)–(3.3) is constrained. We assume that the Hamiltonian $H:\overline{\Omega}\times{\mathbb{R}}^{n}\to{\mathbb{R}}^{n}$ fulfills
$C:=\{(x,p)\in\overline{\Omega}\times{\mathbb{R}}^{n}\ |\ H(x,p)\leq 0\}\quad\text{is closed},$ (3.5)
$C_{x}:=\{p\in{\mathbb{R}}^{n}\ |\ H(x,p)\leq 0\}\quad\text{is convex}\quad\forall x\in\overline{\Omega},$ (3.6)
$\forall(x,p)\in C\ \forall\delta>0\ \exists(x,p_{\delta})\in{C}^{\circ}\ \text{such that}\ |p-p_{\delta}|\leq\delta,$ (3.7)
$\forall(x,p)\in\overline{\Omega}\times{\mathbb{R}}^{n}:\ H(x,p)<0\ \Leftrightarrow\ \exists\lambda\in(0,1):\ \ p/\lambda\in C_{x},$ (3.8)
$\exists 0<\sigma_{*}\leq\sigma^{*}:\quad B_{\sigma_{*}}\subset C_{x}\subset B_{\sigma^{*}}\quad\forall x\in\overline{\Omega}.$ (3.9)
Note that these assumptions are weaker than the corresponding (2.5)–(2.9) and that, in particular, $H$ is not assumed to be continuous.
Due to the definition of viscosity solution (or constrained viscosity solution), in order to solve problem (3.2)–(3.3) we can equivalently replace $H$ by any $\overline{H}$ such that $\{\overline{H}\leq 0\}=\{H\leq 0\}$ and $\{\overline{H}\geq 0\}=\{H\geq 0\}$. This fact is used in the proof of Proposition 3.3 later on.
For all $x\in\overline{\Omega}$ we denote by $\sigma(x,\cdot):{\mathbb{R}}^{n}\to[0,\infty)$ the support function of the nonempty, convex, and closed set $C_{x}$, namely,
$\sigma(x,q)=\sup\{q\cdot p\ |\ H(x,p)\leq 0\}\quad\forall(x,q)\in\overline{\Omega}\times{\mathbb{R}}^{n}.$
We have that $q\mapsto\sigma(x,q)$ is nonnegative and positively $1$-homogeneous. In particular, $0=\sigma(x,0)=\min_{q}\sigma(x,q)$. Assumption (3.9) ensures that
$\sigma_{*}|q|\leq\sigma(x,q)\leq\sigma^{*}|q|\quad\forall(x,q)\in\overline{\Omega}\times{\mathbb{R}}^{n}.$ (3.10)
Let us collect come basic properties of $\sigma$ in the following.
Lemma 3.1 (Properties of $\sigma$).
Under assumptions (3.5)–(3.7) and (3.9), the support function $\sigma$ is continuous and $q\mapsto\sigma(x,q)$ is convex for all $x\in\overline{\Omega}$.
Proof.
For all $x\in\overline{\Omega}$, the function $q\mapsto\sigma(x,q)$ is the support function of the convex set $C_{x}$, hence it is convex.
Let $(x_{j},q_{j})_{j}\in\overline{\Omega}\times{\mathbb{R}}^{n}$ with $(x_{j},q_{j})\to(x,q)$ and, for all $\zeta>0$ let $p_{\zeta}\in C_{x}$ such that $\sigma(x,q)\leq q\cdot p_{\zeta}+\zeta$. Using (3.7) we find $p_{\zeta\delta}$ such that $(x,p_{\zeta\delta})\in C^{\circ}$ and $q\cdot p_{\zeta}\leq q\cdot p_{\zeta\delta}+|q|\delta$. In particular, we have that $(x_{j},p_{\zeta\delta})\in C$ for $j$ large enough, so that $p_{\zeta\delta}\in C_{x_{j}}$, and we conclude that
$\sigma(x,q)\leq q\cdot p_{\zeta}+\zeta\leq q\cdot p_{\zeta\delta}+\zeta+|q|\delta=\lim_{j\to\infty}q_{j}\cdot p_{\zeta\delta}+\zeta+|q|\delta$
$\quad\stackrel{{\scriptstyle p_{\zeta\delta}\in C_{x_{j}}}}{{\leq}}\liminf_{n\to\infty}\sigma(x_{j},q_{j})+\zeta+|q|\delta.$
As $\zeta$ and $\delta$ are arbitrary, this proves that $\sigma$ is lower semicontinuous.
Let us now check upper semicontinuity. For all $j$ we can find $p_{j}\in C_{x_{j}}$ such that $\sigma(x_{j},q_{j})\leq q_{j}\cdot p_{j}+1/j$. Let the subsequence $(j_{k})_{k}$ be such that $\lim_{k\to\infty}\sigma(x_{j_{k}},q_{j_{k}})=\limsup_{j\to\infty}\sigma(x_{j},q_{j})$. As $p_{j_{k}}\in C_{x_{j_{k}}}\subset B_{\sigma^{*}}$ by (3.9), one can extract without relabeling in such a way that $p_{j_{k}}\to p$. As $C$ is closed by (3.5), we have that $(x,p)\in C$, namely, $p\in C_{x}$, and
$\limsup_{n\to\infty}\sigma(x_{j},q_{j})=\lim_{k\to\infty}\sigma(x_{j_{k}},q_{j_{k}})=\lim_{k\to\infty}q_{j_{k}}\cdot p_{j_{k}}=q\cdot p\stackrel{{\scriptstyle p\in C_{x}}}{{\leq}}\sigma(x,q)$
so that $\sigma$ is upper semicontinuous, as well. ∎
In the unconstrained case $\Omega={\mathbb{R}}^{n}$, existence of viscosity solutions to (3.2)–(3.3) is ensured by the following.
Proposition 3.2 (Existence and representation in the unconstrained case).
Let $\Omega={\mathbb{R}}^{n}$. Under assumptions (3.5)–(3.7) and (3.9) problem (3.2)–(3.3) admits a unique nonnegative viscosity solution, which is given by the representation formula
$v(x)=\min\left\{\int_{0}^{1}\sigma(\gamma(s),\gamma^{\prime}(s))\ {{\rm d}s}\ \Big|\ \gamma\in{\rm Lip}_{x}\right\}\quad\forall x\in{\mathbb{R}}^{n}.$ (3.11)
Proof.
We first check that
$C_{x}=\overline{\{p\in{\mathbb{R}}^{n}\ |\ (x,p)\in C^{\circ}\}}=:D_{x}\quad\forall x\in{\mathbb{R}}^{n}.$ (3.12)
The inclusion $D_{x}\subset C_{x}$ is immediate: for all $p\in D_{x}$ we find $p_{j}\to p$ with $H(x,p_{j})\leq 0$. As $C\ni(x,p_{j})\to(x,p)$, the closure of $C$ from (3.5) ensures that $(x,p)\in C$, namely, $p\in C_{x}$.
Let now $p\in C_{x}$. Condition (3.7) implies that we can find a sequence $(p_{\delta})_{\delta}$ with $(x,p_{\delta})\in C^{\circ}$ and $p_{\delta}\to p$ as $\delta\to 0$. This proves that $p\in D_{x}$, hence $C_{x}\subset D_{x}$.
Having proved (3.12), the assertion follows from [44, Thm. 3.15]. ∎
A consequence of Proposition 3.2 is that the minimum in (3.11) is always attained. In the following, we call optimal curve for $x$ a minimizer $\gamma_{x}\in{\rm Lip}_{x}$ in (3.11). Note that optimal curves $\gamma_{x}$ are not necessarily unique. Before moving on, let us mention that the assumptions of Proposition 3.2 could be somehow weakened (see [44]). We work nonetheless in the setting of (3.5)–(3.7) and (3.9), which allows the coupling with the PDE for the activating field $u$.
By dropping the nonnegativity requirement on $v$, uniqueness in Proposition 3.2 fails. An elementary illustration of this fact are the viscosity solutions
$x\mapsto v_{0}(x)=(|x|-1)^{+}\quad\text{and}\quad x\mapsto v_{r}(x)=\min\{v_{0}(x),r-|x|\}\quad\forall r\geq 1$ (3.13)
to problem (3.2)–(3.3) in the unconstrained case $\Omega={\mathbb{R}}^{n}$ for H(x,p)=|p|1H(x,p)=|p|-1 and $V_{0}={B_{1}}$
We now turn to consider the constrained case $\Omega\not={\mathbb{R}}^{n}$ under assumption (2.1). In contrast to the unconstrained case, uniqueness of viscosity solutions to the constrained case cannot be restored by simply requiring nonnegativity: Consider again the case of (3.13) and take $\Omega=B_{3}$. The restrictions to $\Omega$ of all functions in (3.13) are viscosity solutions to the constrained problem. Nonetheless, we have that $v_{0}$, $v_{r}\geq 0$ in $\Omega$ for all $r\geq 3$, so that nonnegativity alone is not selective.
Aiming at identifying some possible selection principle among viscosity solutions, we are interested in a constrained version of the representation formula (3.11). The existence result in the constrained case reads as follows.
Proposition 3.3 (Existence and representation in the constrained case).
Assume (2.1)–(2.2), (3.1), and (3.5)–(3.9). Then, $v\in C(\overline{\Omega})$ given by
$v(x)=\min\Bigg\{\int_{0}^{1}\sigma(\gamma(s),\gamma^{\prime}(s))\ {{\rm d}s}\ \Big|\$$\gamma\in{\rm Lip}_{x},\ \gamma(\cdot)\in\overline{\Omega}\Bigg\}\quad\forall x\in\overline{\Omega}.$ (3.14)
is a viscosity solution to problem (3.2)–(3.3). For $H$ continuous, under the following uniform but one-sided version of (3.8)
$\forall\delta>0\ \exists\lambda\in(0,1)\ \forall(x,p)\in\overline{\Omega}\times{\mathbb{R}}^{n}:\ H(x,p)<-\delta\ \Rightarrow\ p/\lambda\in C_{x},$ (3.15)
the function $v$ is a constrained viscosity solution.
Remark 3.4 (Alternative assumptions).
For $H$ continuous, assumption (3.5) is of course satisfied. If in addition $p\in{\mathbb{R}}^{n}\mapsto H(x,p)$ is convex and $\inf H(x,\cdot)<0$ for all $x\in\overline{\Omega}$, property (3.7) follows. Moreover, if $\Omega$ is convex, assumption (3.8) can be dropped. This is also the case if $\overline{\Omega}$ is homeomorphic to a convex and closed set $D\subset{\mathbb{R}}^{n}$ via a global homeomorphism $h:{\mathbb{R}}^{n}\to{\mathbb{R}}^{n}$, namely, $\overline{\Omega}=h(D)$, see Remark 3.6 below.
Remark 3.5 (Uniqueness).
Even in the more regular case of (3.15), if $p\mapsto H(x,p)$ is not convex, no uniqueness can be expected in general, see [40]. On the other hand, if $H$ is Lipschitz continuous and $p\mapsto H(x,p)$ is convex and $H(x,0)<0$ for all $x\in\overline{\Omega}$, [14, Prop. IX.3] guarantees that $v$ given by (3.14) is the unique nonnegative viscosity solution to problem (3.2)–(3.3). This in particular applies to the case of the generalized eikonal equation $H(x,p)=\gamma(x)|p|-1$ for $\gamma$ Lipschitz continuous, bounded, and uniformly positive.
Proof of Proposition 3.3.
The proof relies on a penalization procedure in the spirit of [14, Sec. VII], and is divided into steps. At first, we reformulate the problem by equivalently replacing $H$ by another Hamiltonian $\overline{H}$ having $\{\overline{H}\leq 0\}=\{H\leq 0\}$ and $\{\overline{H}\geq 0\}=\{H\geq 0\}$ (Step 1). Then, we provide extensions $\overline{H}_{\varepsilon}$ of $\overline{H}$ to all $x\in{\mathbb{R}}^{n}$, depending on a small parameter $\varepsilon>0$ (Step 2). For all such $\overline{H}_{\varepsilon}$ we solve the unconstrained problem in ${\mathbb{R}}^{n}$ by the representation formula (3.11). This defines functions $v_{\varepsilon}\in C({\mathbb{R}}^{n})$ (Step 3). By taking $\varepsilon\to 0$, we prove that $v_{\varepsilon}\to v$ locally uniformly in $\overline{\Omega}$, that $v\in C(\overline{\Omega})$ is a viscosity solution to (3.2)–(3.3) (Step 4), and that the representation formula (3.14) holds (Step 5). Eventually, under (3.15) we prove that $v$ is a constrained viscosity solution (Step 6).
Step 1: Replacing $H$ by $\overline{H}$. Define
$\overline{H}(x,p):=\inf\{\lambda>0\ |\ p/\lambda\in C_{x}\}-1\quad\forall(x,p)\in\overline{\Omega}\times{\mathbb{R}}^{n}.$ (3.16)
For all $x\in\overline{\Omega}$, $\overline{H}(x,\cdot)$ is the standard Minkowski functional of the convex set $C_{x}$, up to an additive constant. As $0\in B_{\sigma_{*}}\subset C_{x}$ we have that $\overline{H}(x,0)=-1$ and $p\mapsto\overline{H}(x,p)$ is positively $1$-homogeneous. Moreover, for all $(x,p)\in\overline{\Omega}\times{\mathbb{R}}^{n}$ with $H(x,p)\leq 0$ one has that $(x,p/1)\in C$, hence $\overline{H}(x,p)\leq 0$. On the other hand, if $\overline{H}(x,p)\leq 0$ there exists $\lambda_{\delta}\to 1$ such that $(x,p/\lambda_{\delta})\in C$ and the closure of $C$ from (3.5) ensures that $(x,p)\in C$, namely, $H(x,p)\leq 0$. This proves that $\{\overline{H}\leq 0\}=\{H\leq 0\}$. As assumption (3.8) directly implies that $\{H<0\}=\{\overline{H}<0\}$, by taking the complements we have that $\{\overline{H}\geq 0\}=\{H\geq 0\}$. We are hence allowed to replace $H$ by $\overline{H}$ in (3.2)–(3.3) without changing the corresponding viscosity solutions. Note more precisely that $C_{x}=\{p\in{\mathbb{R}}^{n}\ |\ \overline{H}(x,p)\leq 0\}$ for all $x\in\overline{\Omega}$, so that $\sigma$ and formula (3.14) remains unchanged.
Step 2: Extension of $\overline{H}$. For all $\varepsilon\in(0,\min\{\sqrt{\overline{\varepsilon}},1/\sigma^{*}\})$ we define
$C_{x\varepsilon}:=\left(1-\frac{d(x,\Omega)}{\varepsilon^{2}}\right)^{+}C_{\pi(x)}+\min\left\{\frac{d(x,\Omega)}{\varepsilon^{2}},1\right\}\overline{B_{1/\varepsilon}}\quad\forall x\in{\mathbb{R}}^{n}.$
In the above right-hand side, the first set is empty unless $d(x,\Omega)<\varepsilon^{2}$. In this case, as $\varepsilon^{2}<\overline{\varepsilon}$ we have that $d(x,\Omega)<\overline{\varepsilon}$ and the projection $\pi(x)={\rm arg\,min}\{|x-\omega|\>|\>\omega\in\overline{\Omega}\}$ is well defined. Specifically, we have $C_{x\varepsilon}=C_{x}$ for all $x\in\overline{\Omega}$ and $C_{x\varepsilon}=\overline{B_{1/\varepsilon}}$ if $d(x,\Omega)\geq\varepsilon^{2}$.
Let us now define $\overline{H}_{\varepsilon}:{\mathbb{R}}^{n}\times{\mathbb{R}}^{n}\to[0,\infty)$ as
$\overline{H}_{\varepsilon}(x,p)=\inf\{\lambda>0\ |\ p/\lambda\in C_{x\varepsilon}\}-1$
and observe that $C_{x\varepsilon}=\{p\in{\mathbb{R}}^{n}\>|\>\overline{H}_{\varepsilon}(x,p)\leq 0\}$ for all $x\in{\mathbb{R}}^{n}$.
In the remainder of this step, we check that assumptions (3.5)–(3.9) for $H$ in $\overline{\Omega}$ imply that $\overline{H}_{\varepsilon}$ fulfills assumptions (3.5)–(3.7) and (3.9) for $\Omega={\mathbb{R}}^{n}$. To simplify notation, in the following we use $\alpha(x):=(1-d(x,\Omega)/\varepsilon^{2})^{+}$ and $\beta(x):=\min\{d(x,\Omega)/\varepsilon^{2},1\}$ for all $x\in{\mathbb{R}}^{n}$.
At first, let us check that
$C_{\varepsilon}:=\{(x,p)\in{\mathbb{R}}^{n}\times{\mathbb{R}}^{n}\ |\ \overline{H}_{\varepsilon}(x,p)\leq 0\}$
is closed. Take a sequence $(x_{j},p_{j})_{j}\in C_{\varepsilon}$ with $(x_{j},p_{j})\to(x,p)$. Then, we can find $q_{j}\in C_{\pi(x_{j})}$ and $r_{j}\in\overline{B_{1/\varepsilon}}$ such that $p_{j}=\alpha(x_{j})q_{j}+\beta(x_{j})r_{j}$. Upon possibly extracting not relabeled subsequences and using (3.5) and (3.9) we can assume that $q_{j}\to q\in C_{\pi(x)}$ and $r_{j}\to r\in\overline{B_{1/\varepsilon}}$. Hence, $p_{j}\to\alpha(x)q+\beta(x)r\in C_{x\varepsilon}$. This proves that $(x,p)\in C_{\varepsilon}$. In particular, $C_{\varepsilon}$ is closed and (3.5) holds for $\overline{H}_{\varepsilon}$ on the whole ${\mathbb{R}}^{n}$.
For all $x\in{\mathbb{R}}^{n}$, $C_{x\varepsilon}$ is the sum of two convex sets and hence convex. This implies that (3.6) holds for $\overline{H}_{\varepsilon}$ for all $x\in{\mathbb{R}}^{n}$.
In order to check (3.7) for $\overline{H}_{\varepsilon}$ we fix $(x,p)\in C_{\varepsilon}$ and distinguish some cases:
  • Case $x\in\Omega$. From $p\in C_{x\varepsilon}=C_{x}$, for all $\delta>0$ we use (3.7) for $H$ in order to find $p_{\delta}$ with $|p-p_{\delta}|\leq\delta$ and $(x,p_{\delta})\in C^{\circ}$. As $\Omega$ is open, we directly get that $(x,p_{\delta})\in C_{\varepsilon}^{\circ}$, as well.
[…]
    $\alpha^{-1}(\widetilde{x})\widetilde{p}$ $=p_{\delta}+s+(\alpha^{-1}(\widetilde{x}){-}1)\widetilde{p}=p_{\delta}+s+(\alpha^{-1}(\widetilde{x}){-}1)(p_{\delta}+s)$
[…]
    $=:\alpha(\widetilde{x})\widehat{q}+\beta(\widetilde{x})\widehat{r}$.
    For $y$ and $s$ small, $\widehat{q}$ and $\widehat{r}$ are arbitrarily close to $q_{\delta}$ and $r_{\delta}$, respectively. In particular, $\widehat{p}\in C_{\pi(\widetilde{x})}$ since $(\pi(x),p_{\delta})\in C^{\circ}$ and $\widehat{r}\in\overline{B_{1/\varepsilon}}$. This proves that $\widetilde{p}\in C_{\widetilde{x}\varepsilon}$ or, equivalently, $(\widetilde{x},\widetilde{p})\in C_{\varepsilon}$. As a consequence, we have that $(x,p_{\delta})\in C_{\varepsilon}^{\circ}$.
[…]
    $\beta^{-1}(\widetilde{x})\widetilde{p}$ $=p_{\delta}+s+(\beta^{-1}(\widetilde{x}){-}1)\widetilde{p}=p_{\delta}+s+(\beta^{-1}(\widetilde{x}){-}1)(p_{\delta}+s).$
    As $y$ and $s$ are small, we have that $\beta^{-1}(\widetilde{x})\widetilde{p}\in\overline{B_{1/\varepsilon}}$ and we have proved that $\widetilde{p}=\beta(\widetilde{x})(\beta^{-1}(\widetilde{x})\widetilde{p})\in C_{\widetilde{x}\varepsilon}$. This implies that $(\widetilde{x},\widetilde{p})\in C_{\varepsilon}$, hence $(x,p_{\delta})\in C^{\circ}_{\varepsilon}$.
  • Case $d(x,\Omega)>\varepsilon^{2}$. We have that $C_{\widetilde{x}\varepsilon}=\overline{B_{1/\varepsilon}}$ for all $\widetilde{x}$ close to $x$. For all $\delta>0$, choosing $p_{\delta}\in B_{1/\varepsilon}$ with $|p-p_{\delta}|\leq\delta$ we have $(x,p_{\delta})\in C^{\circ}_{\varepsilon}$.
As $\varepsilon<1/\sigma^{*}<1/\sigma_{*}$, we can use (3.9) and deduce for all $x\in{\mathbb{R}}^{n}$ that
$B_{\sigma_{*}}=(\alpha(x)+\beta(x))B_{\sigma_{*}}\subset\alpha(x)B_{\sigma_{*}}+\beta(x)B_{1/\varepsilon}$
$\quad\subset C_{x\varepsilon}\subset\alpha(x)B_{\sigma^{*}}+\beta(x)\overline{B_{1/\varepsilon}}\subset(\alpha(x)+\beta(x))\overline{B_{1/\varepsilon}}=\overline{B_{1/\varepsilon}}.$
This proves that $\overline{H}_{\varepsilon}$ fulfills (3.9) in the whole ${\mathbb{R}}^{n}$, but with $\sigma^{*}$ replaced by $B_{r}$ for $r>1/\varepsilon$. In particular,
$B_{\sigma_{*}}\subset C_{x\varepsilon}\subset{B_{2/\varepsilon}}\quad\forall x\in{\mathbb{R}}^{n}.$ (3.17)
Step 3: Solution of the approximating unconstrained problem. Since $\overline{H}_{\varepsilon}$ fulfills assumptions (3.5)–(3.7) and (3.9) in ${\mathbb{R}}^{n}$ (with $\sigma^{*}$ replaced by $2/\varepsilon$), Proposition 3.2 applies and the function
$v_{\varepsilon}(x)=\min\Bigg\{\int_{0}^{1}\sigma_{\varepsilon}(\gamma(s),\gamma^{\prime}(s))\ {{\rm d}s}\ \Big|\ \gamma\in{\rm Lip}_{x}\Bigg\}\quad\forall x\in{\mathbb{R}}^{n}$ (3.18)
is the unique nonnegative viscosity solution to (3.2)–(3.3) for $\Omega={\mathbb{R}}^{n}$ and $H$ replaced by $\overline{H}_{\varepsilon}$. In (3.18), the support function $\sigma_{\varepsilon}$ is defined as
$\sigma_{\varepsilon}(x,q)$ $:=\sup\{q\cdot p\ |\ \overline{H}_{\varepsilon}(x,p)\leq 0\}\quad\forall(x,q)\in{\mathbb{R}}^{n}\times{\mathbb{R}}^{n}$
Owing to (3.17) we have that
$\sigma_{*}|q|\leq\sigma_{\varepsilon}(x,q)\leq\frac{2}{\varepsilon}|q|\quad\forall(x,q)\in{\mathbb{R}}^{n}\times{\mathbb{R}}^{n}.$ (3.19)
$\sigma_{\varepsilon}(x,q)=\sigma(x,q)\leq\sigma^{*}|q|\quad\forall(x,q)\in\overline{\Omega}\times{\mathbb{R}}^{n}.$ (3.20)
Moreover, as $C_{\pi(x)}=(\alpha(x)+\beta(x))C_{\pi(x)}\subset\alpha(x)C_{\pi(x)}+\beta(x)\overline{B_{1/\varepsilon}}=C_{x\varepsilon}$ for all $x\in{\mathbb{R}}^{n}$ such that $d(x,\Omega)\leq\varepsilon^{2}$, it follows that
$\sigma(\pi(x),q)\leq\sigma_{\varepsilon}(x,q)\quad\forall(x,q)\in{\mathbb{R}}^{n}\times{\mathbb{R}}^{n}\ \ \text{with}\ \ d(x,\Omega)\leq\varepsilon^{2}.$ (3.21)
Step 4: Estimate and limit for $\varepsilon\to 0$. For all $x,\,y\in\overline{\Omega}$, let $\gamma_{y}\in{\rm Lip}_{y}$ be optimal for $v_{\varepsilon}$ at $y$ and $\gamma\in W^{1,\infty}(0,1;\overline{\Omega})$ be a geodesic from $y$ to $x$. By considering the concatenated curve $\widetilde{\gamma}\in{\rm Lip}_{x}$ given by
$\widetilde{\gamma}(s)=\left\{\begin{array}[]{ll}\gamma_{y}(2s)\quad&\text{for}\ s\in[0,1/2],\\[2.84526pt] \gamma(2s-1)&\text{for}\ s\in(1/2,1],\end{array}\right$.
we can bound
$v_{\varepsilon}(x)-v_{\varepsilon}(y)\leq\int_{0}^{1}\sigma_{\varepsilon}(\widetilde{\gamma}(s),\widetilde{\gamma}^{\prime}(s))\,{\rm d}s-\int_{0}^{1}\sigma_{\varepsilon}(\gamma_{y}(s),\gamma^{\prime}_{y}(s))\,{\rm d}s$
[…]
By exchanging the roles of $x$ and $y$, this proves that $|\nabla v_{\varepsilon}|\leq\sigma^{*}L$ almost everywhere in $\Omega$. A diagonal extraction argument allows to find $v\in C(\overline{\Omega})$ and a not relabeled subsequence $v_{\varepsilon}$ such that
$v_{\varepsilon}\to v\quad\text{locally uniformly in}\ \overline{\Omega}.$
It is a standard matter to prove that $v$ is a viscosity solution to the constrained problem (3.2)–(3.3). Indeed, for all $x\in\Omega\setminus\overline{V_{0}}$ and all $\varphi\in C^{1}({\mathbb{R}}^{n})$ with $\varphi(x)-v(x)=\min(\varphi-v)$ relative to $\Omega$ we use [19, Prop. 4.3] in order to find $\Omega\setminus\overline{V_{0}}\ni x_{\varepsilon}\to x$ and $\varphi_{\varepsilon}\in C^{1}({\mathbb{R}}^{n})$ with $\varphi_{\varepsilon}(x_{\varepsilon})-v_{\varepsilon}(x_{\varepsilon})=\min(\varphi_{\varepsilon}-v_{\varepsilon})$ relative to $\Omega$ and $\nabla\varphi(x_{\varepsilon})\to\nabla\varphi(x)$. As $\overline{H}(z,\cdot)=\overline{H}_{\varepsilon}(z,\cdot)$ for $z\in\Omega\setminus\overline{V_{0}}$ one gets $0\geq\lim_{\varepsilon\to 0}\overline{H}_{\varepsilon}(x_{\varepsilon},\nabla\varphi(x_{\varepsilon}))=\overline{H}(x,\nabla\varphi(x))$, proving that $v$ is a viscosity subsolution to (3.2) in $\Omega\setminus\overline{V_{0}}$. Similarly one proves that $v$ is a viscosity supersolution, hence a viscosity solution.
Step 5: Representation formula. We now prove that the limit $v\in C(\overline{\Omega})$ is represented as in (3.14). Fix $x\in\overline{\Omega}$. If $x\in\overline{V}_{0}$ the formula trivially holds, so that we can assume $x\in\overline{\Omega}\setminus\overline{V_{0}}$. Let $\gamma\in{\rm Lip}_{x}$ with $\gamma(s)\in\overline{\Omega}$ for all $s\in[0,1]$. Then,
$v(x)=\lim_{\varepsilon\to 0}v_{\varepsilon}(x)\leq\int_{0}^{1}\sigma_{\varepsilon}(\gamma(s),\gamma^{\prime}(s))\,{\rm d}s\stackrel{{\scriptstyle\gamma(\cdot)\in\overline{\Omega}}}{{=}}\int_{0}^{1}\sigma(\gamma(s),\gamma^{\prime}(s))\,{\rm d}s$ (3.23)
so that the right-hand side of (3.14) is an upper bound for $v(x)$.
We are left with checking that the minimum in (3.14) is actually attained, namely, that there exists an optimal curve in $\overline{\Omega}$ realizing the equality in (3.23). To this aim, let $\gamma_{x\varepsilon}\in{\rm Lip}_{x}$ be optimal curves for $x$ under $\sigma_{\varepsilon}$ and note that the corresponding lengths $\ell_{\varepsilon}$ are uniformly bounded as
[…]
and $v_{\varepsilon}(x)\to v(x)$. In particular, $\ell_{\rm max}:=\sup_{\varepsilon}\ell_{\varepsilon}<\infty$. On the other hand, the lengths $\ell_{\varepsilon}$ are bounded below by $d(x,\overline{V}_{0})>0$. We indicate by $\widetilde{\gamma}_{x\varepsilon}\in W^{1,\infty}(0,\ell_{\varepsilon};{\mathbb{R}}^{n})$ the arc-length reparametrizations of $\gamma_{x\varepsilon}$, see [2, Lemma 1.1.4(b), p. 25], and trivially extend them to $0,\ell_{\rm\max}]$ without changing notation. As the curves $\widetilde{\gamma}_{x\varepsilon}$ are uniformly Lipschitz continuous and $\widetilde{\gamma}_{x\varepsilon}(\ell_{\rm max})=x$, one can extract not relabeled subsequences in such a way that $\ell_{\varepsilon}\to\ell\in[d(x,\overline{V}_{0}),\ell_{\rm max}]$ and $\widetilde{\gamma}_{x\varepsilon}\to\widetilde{\gamma}_{x}$ uniformly in $0,\ell_{\rm max}]$ and weakly* in $W^{1,\infty}(0,\ell_{\rm max};{\mathbb{R}}^{n})$. Note that $|\widetilde{\gamma}_{x}^{\prime}|\leq 1$ a.e. in $(0,\ell)$ and $\widetilde{\gamma}_{x}^{\prime}=0$ a.e. in $(\ell,\ell_{\rm max})$. Hence, $\widetilde{\gamma}_{x}(\ell)=\widetilde{\gamma}_{x}(\ell_{\rm max})=x$, and $\overline{V_{0}}\ni\widetilde{\gamma}_{x\varepsilon}(0)\to\widetilde{\gamma}_{x}(0)$. As $\overline{V_{0}}$ is closed, the latter convergence ensures that $\widetilde{\gamma}_{x}(0)\in\overline{V_{0}}$, as well.
Let us now check that $\widetilde{\gamma}_{x}(s)\in\overline{\Omega}$ for all $s\in[0,\ell]$. To this aim, define
$A_{\varepsilon}:=\{s\in[0,\ell_{\varepsilon}]\ |\ d(\widetilde{\gamma}_{x\varepsilon}(s),\overline{\Omega})\geq\varepsilon^{2}\}.$
For almost all $s\in A_{\varepsilon}$ we have that
$\sigma_{\varepsilon}(\widetilde{\gamma}_{x\varepsilon}(s),\widetilde{\gamma}_{x\varepsilon}^{\prime}(s))=\frac{1}{\varepsilon}|\widetilde{\gamma}_{x\varepsilon}^{\prime}(s)|=\frac{1}{\varepsilon}.$
This allows to estimate $|A_{\varepsilon}|$ as follows
$\frac{1}{\varepsilon}|A_{\varepsilon}|=\int_{A_{\varepsilon}}\sigma_{\varepsilon}(\widetilde{\gamma}_{x\varepsilon}(s),\widetilde{\gamma}_{x\varepsilon}^{\prime}(s))\,{\rm d}s\leq\int_{0}^{\ell_{\varepsilon}}\sigma_{\varepsilon}(\widetilde{\gamma}_{x\varepsilon}(s),\widetilde{\gamma}_{x\varepsilon}^{\prime}(s))\,{\rm d}s=v_{\varepsilon}(x)\to v(x),$
implying that
$|A_{\varepsilon}|\leq c\varepsilon$ (3.25)
for some $c>0$ depending on $v(x)$ only. Hence, the length of the portion of the arc-parametrized curve $\widetilde{\gamma}_{x\varepsilon}$ laying in the complement of $\overline{\Omega}+{B_{\varepsilon^{2}}}$ is smaller than $c\varepsilon$. This in particular entails that
$d(\widetilde{\gamma}_{x\varepsilon}(s),\overline{\Omega}+{B_{\varepsilon^{2}}})\leq c\varepsilon\quad\forall s\in[0,\ell_{\varepsilon}]$
and by the triangle inequality we get
$d(\widetilde{\gamma}_{x\varepsilon}(s),\overline{\Omega})\leq d(\widetilde{\gamma}_{x\varepsilon}(s),\overline{\Omega}+{B_{\varepsilon^{2}}})+\varepsilon^{2}\leq c\varepsilon+\varepsilon^{2}\quad\forall s\in[0,\ell_{\varepsilon}].$ (3.26)
By passing to the limit as $\varepsilon\to 0$ we obtain that $\widetilde{\gamma}_{x}(s)\in\overline{\Omega}$ for all $s\in[0,\ell]$.
Let now $\delta\in(0,\ell)$ be given and choose $\varepsilon_{j}:=2^{-j}\delta$ for $j\geq 1$. For $j$ large enough, one has that $c\varepsilon_{j}+\varepsilon^{2}_{j}<\overline{\varepsilon}$. Define $A:=\cup_{j}A_{\varepsilon_{j}}$ so that $|A|\leq\sum_{j}|A_{\varepsilon_{j}}|\leq c\delta$. By using (3.21) we find that
$v(x)$ $=\lim_{j\to\infty}v_{\varepsilon_{j}}(x)=\lim_{j\to\infty}\int_{0}^{\ell_{\varepsilon_{j}}}\sigma_{\varepsilon_{j}}(\widetilde{\gamma}_{x\varepsilon_{j}}(s),\widetilde{\gamma}_{x\varepsilon_{j}}^{\prime}(s))\,{\rm d}s$
$\geq\liminf_{j\to\infty}\int_{[0,\ell_{\varepsilon_{j}}]\setminus A_{\varepsilon_{j}}}\sigma_{\varepsilon_{j}}(\widetilde{\gamma}_{x\varepsilon_{j}}(s),\widetilde{\gamma}_{x\varepsilon_{j}}^{\prime}(s))\,{\rm d}s$
[…]
$\geq\liminf_{j\to\infty}\int_{[0,\ell-\delta]\setminus A}\sigma(\pi(\widetilde{\gamma}_{x\varepsilon_{j}}(s)),\widetilde{\gamma}_{x\varepsilon_{j}}^{\prime}(s))\,{\rm d}s$
$\geq\int_{[0,\ell-\delta]\setminus A}\sigma(\pi(\widetilde{\gamma}_{x}(s)),\widetilde{\gamma}_{x}^{\prime}(s))\,{\rm d}s$
$\geq\int_{0}^{\ell}\sigma(\pi(\widetilde{\gamma}_{x}(s)),\widetilde{\gamma}_{x}^{\prime}(s))\,{\rm d}s-\sigma^{*}\delta-\sigma^{*}|A|$
$\hskip-10.243pt\stackrel{{\scriptstyle\widetilde{\gamma}_{x}(\cdot)\in\overline{\Omega}}}{{\geq}}\int_{0}^{\ell}\sigma(\widetilde{\gamma}_{x}(s),\widetilde{\gamma}_{x}^{\prime}(s))\,{\rm d}s-\sigma^{*}(1+c)\delta$
where we have used the continuity of $\sigma$ and $\pi$, the convexity of $\sigma(x,\cdot)$, see Lemma 3.1, and the classical [20, Thm. 3.23, p. 96]. As $\delta$ is arbitrary, by defining $\gamma_{x}\in{\rm Lip}_{x}$ as $\gamma_{x}(s)=\widetilde{\gamma}_{x}(\ell s)$ for all $s\in[0,1]$, this proves that
$v(x)\geq\int_{0}^{\ell}\sigma(\widetilde{\gamma}_{x}(s),\widetilde{\gamma}_{x}^{\prime}(s))\,{\rm d}s=\int_{0}^{1}\sigma(\gamma_{x}(s),\gamma_{x}^{\prime}(s))\,{\rm d}s.$
Together with the upper bound (3.23), we have checked that $\gamma_{x}$ is optimal for $x$ and the representation formula (3.14) holds.
Step 6: Boundary behavior. Let now $H$ be continuous and let (3.15) hold. Assume by contradiction that $v$ is not a constrained viscosity solution. In particular, assume that there exist $x\in\partial\Omega\setminus\overline{V_{0}}$ and $\varphi\in C^{1}({\mathbb{R}}^{n})$ such that $\varphi-v$ has a local maximum at $x$ but $H(x,\nabla\varphi(x))<0$. Let $\gamma\in{\rm Lip}_{x}$ realize the minimum in (3.14) and assume with no loss of generality that $|\gamma^{\prime}|=\ell>0$ a.e. For all $r>0$ we can find $t\in(0,1)$ such that $\gamma(s)\in B_{r}(x)$ for all $s\in[t,1]$. Owing to the continuity of $H$ and $\nabla\varphi$ and possibly choosing a larger $t$, we find $\delta>0$ such that $H(\gamma(s),\nabla\varphi(\gamma(s))<-\delta$ for all $s\in[t,1]$. By (3.15) there exists $\lambda\in(0,1)$ with $H(\gamma(s),\nabla\varphi(\gamma(s))/\lambda)\leq 0$ for all $s\in[t,1]$. Using the very definition of $\sigma$ this entails that
$\gamma^{\prime}(s)\cdot\nabla\varphi(\gamma(s))\leq\lambda\sigma(\gamma(s),\gamma^{\prime}(s))\quad\text{for a.e.}\ s\in(t,1).$ (3.27)
We can now compute
[…]
$\quad=\int_{t}^{1}\sigma(\gamma(s),\gamma^{\prime}(s))\,{\rm d}s+(\lambda-1)\int_{t}^{1}\sigma(\gamma(s),\gamma^{\prime}(s))\,{\rm d}s$
$\quad\leq\int_{0}^{1}\sigma(\gamma(s),\gamma^{\prime}(s))\,{\rm d}s-\int_{0}^{t}\sigma(\gamma(s),\gamma^{\prime}(s))\,{\rm d}s+(\lambda-1)(1-t)\ell\sigma_{*}$
$\quad\leq v(x)-v(\gamma(t))+(\lambda-1)(1-t)\ell\sigma_{*}<v(x)-v(\gamma(t)).$ (3.28)
This implies that $\varphi(\gamma(t))-v(\gamma(t))>\varphi(x)-v(x)$. As $r>0$ is arbitrary, we have proved that $x$ is not a local maximum of $\varphi-v$, which is a contradiction. ∎
Remark 3.6 (Alternative assumptions, continued).
In case $\Omega$ is convex the proof of Proposition 3.3 can be simplified by skipping Step 1 and arguing directly on the extension $H_{\varepsilon}(x,p)=H(\pi(x),b_{\varepsilon}(x)p)$ where $\pi$ is the projection on $\overline{\Omega}$ and $b_{\varepsilon}(x)=a_{\varepsilon}(d(x,\overline{\Omega})/\varepsilon^{2})$ for some $a_{\varepsilon}:{\mathbb{R}}\to[\varepsilon,1]$ smooth with $a_{\varepsilon}(r)=1$ for $r\leq 1$ and $a_{\varepsilon}(r)=\varepsilon$ for $r\geq 2$. In fact, in this case one only needs $\pi$ to be a continuous map from ${\mathbb{R}}^{n}$ to $\overline{\Omega}$ coinciding with the identity on $\overline{\Omega}$. Hence, one could consider also some nonconvex $\overline{\Omega}=h(\overline{D})$ as indicated in Remark 3.4 by replacing $\pi$ in $H$ by $h\cdot\pi_{D}\circ h^{-1}$, where now $\pi_{D}$ is the projection on $D$.
Let us record some basic estimates from the proof of Proposition 3.3 in the following.
Proposition 3.7 (Estimates on $v$).
Under the assumptions of Proposition 3.3, let $v$ be given by (3.14). Then, the length $\ell_{x}$ of any optimal curve $\gamma_{x}\in{\rm Lip}_{x}$ with $\gamma_{x}(\cdot)\in\overline{\Omega}$ fulfills
$\ell_{x}\leq\frac{v(x)}{\sigma_{*}}\quad\forall x\in\overline{\Omega}.$ (3.29)
Moreover, we have that
$|\nabla v|\leq\sigma^{*}L\quad\text{a.e. in}\ \Omega,$ (3.30)
$\forall R>0:\ v(x)\leq c_{1}(R)\quad\forall x\in\overline{\Omega}\cap B_{R},$ (3.31)
where $c_{1}(R):=\sigma^{*}L(|x_{0}|+R)$ for all $R>0$.
Proof.
The bound (3.29) on the length $\ell_{x}$ follows by passing to the limit in (3.24) as $\varepsilon\to 0$. The Lipschitz bound (3.30) can be obtained by repeating the argument leading to (3.22) for the function $v$, using (3.14).
For (3.31), fix any $x\in\overline{\Omega}\cap B_{R}$ and let $\gamma\in{\rm Lip}_{x}$ be a geodesic in $\overline{\Omega}$ from $x_{0}$ to $x$. We have
[…]
In view of the coupling with the elliptic problem for the activation field $u$ in Section 4, the regularity of the sublevels $V(t)$ plays a crucial role. We have the following.
Proposition 3.8 (John regularity of $V(t)$).
Under the assumptions of Proposition 3.3, assume additionally (2.3) and that $v$ is given by (3.14). Then, for all $T>0$ there exists $\kappa(T)\in(0,1]$ such that $V(t)=\{x\in\overline{\Omega}\ |\ v(x)<t\}$ are John domains with respect to $x_{0}$ with John constant $\kappa(T)$ for all $t\in[0,T]$, where $\kappa(T)$ depends on on $\sigma_{*}$, $\sigma^{*}$, $\overline{\varepsilon}$, $\|\nabla\eta\|_{L^{\infty}}$, $\kappa_{0}$, $d(x_{0},\partial\Omega)$ (although not indicated in the notation), and $T$, and $\kappa(T)\to 0$ as $T\to\infty$. Moreover, we have that
$d_{H}(V(t),V(s))\leq\frac{1}{\sigma_{*}}(t-s)\quad\forall 0<s\leq t.$ (3.32)
Finally, for all $T>0$ there exists $\mu(T)\in(0,1)$ and $c_{H}(T)>0$ such that
$|V(t)\setminus V(s)|\leq c_{H}(T)(t-s)^{\mu(T)}\quad\forall 0<s\leq t\leq T,$ (3.33)
where $\mu(T)$ and $c_{H}(T)$ have the same dependencies of $\kappa(T)$.
In case $\Omega={\mathbb{R}}^{n}$, the John regularity of $V(t)$ is proved in [22, Thm. 1.1]. In such unconstrained case, the assumptions on $H$ can be weakened and the constant $\kappa$ turns out to be depending just on $\kappa_{0}$ and, in particular, is independent of $T$. The constrained case $\Omega\not={\mathbb{R}}^{n}$ is more delicate, as the presence and the regularity of the boundary $\partial\Omega$ plays a role. Note nonetheless that John regularity is sharp for problem (3.2)–(3.3), see [22] for counterexamples.
Proof of Proposition 3.8.
We divide the proof into steps. At first, we find an arc-length parametrized curve $\gamma$ in $\overline{\Omega}$ such that $d(\gamma(s),\partial V(t)\setminus\partial\Omega)\geq\overline{\kappa}s$ for some $\overline{\kappa}>0$ (Step 1, Figure 2). Then, starting from $\gamma$ we find a second curve $\widetilde{\gamma}$ such that $d(\widetilde{\gamma}(s),\partial V(t))\geq\widetilde{\kappa}s$ for some $0<\widetilde{\kappa}\leq\overline{\kappa}$ (Step 2, Figure 3). Eventually, we reparametrize $\widetilde{\gamma}$ by arc length and conclude that $V(t)$ is John domain (Step 3). The Lipschitz continuity of the set-valued map $t\mapsto V(t)$ with respect to the Hausdorff distance and the Hölder continuity with respect to the $L^{1}$ distance are proved in Step 4 and Step 5, respectively.
Step 1: The curve $\gamma$. For all $T>0$ fixed, let $t\in(0,T]$ and $x\in V(t)$. Let $\widehat{\gamma}_{1}\in{\rm Lip}_{x}$ with $\widehat{\gamma}_{1}(\cdot)\in\overline{\Omega}$ be optimal for $x$, indicate by $\gamma_{1}:[0,\ell_{\gamma_{1}}]\to\overline{\Omega}$ its arc-length reparametrization (see, e.g., [2, Lemma 1.1.4(b), p. 25]), and set $y=\gamma_{1}(0)\in\overline{V_{0}}$. Moreover, using the fact that $\overline{V_{0}}$ is a John domain, set $\gamma_{2}:[0,\ell_{\gamma_{2}}]\to\overline{V_{0}}$ to be such that $\gamma_{2}(0)=y$, $\gamma_{2}(\ell_{\gamma_{2}})=x_{0}$, $|\gamma_{2}^{\prime}|=1$ a.e., and $d(\gamma_{2}(s),\partial V_{0})\geq\kappa_{0}s$ for all $s\in[0,\ell_{\gamma_{2}}]$. Set $\ell_{\gamma}=\ell_{\gamma_{1}}+\ell_{\gamma_{2}}$ and define $\gamma:[0,\ell_{\gamma}]\to\overline{\Omega}$ by
$\gamma(s):=\left\{\begin{array}[]{ll}\gamma_{1}(\ell_{\gamma_{1}}-s)\quad&\text{for}\ s\in[0,\ell_{\gamma_{1}}],\\ \gamma_{2}(s-\ell_{\gamma_{1}})\quad&\text{for}\ s\in(\ell_{\gamma_{1}},\ell_{\gamma}],\end{array}\right$.
see Figure 2. In particular, $\gamma(0)=x$, $\gamma(\ell_{\gamma})=x_{0}$, and $|\gamma^{\prime}|=1$ a.e. We devote the rest of this step to prove that, setting
$\overline{\kappa}=\frac{\sigma_{*}}{2\sigma^{*}+\sigma_{*}}\kappa_{0},$ (3.34)
one has that
$d(\gamma(s),\partial V(t)\setminus\partial\Omega)\geq\overline{\kappa}s\quad\forall s\in[0,\ell_{\gamma}].$ (3.35)
This follows from adapting the argument of [22, Thm. 1.1]. We argue by proving (3.35) separately on the three intervals $0,\ell_{\gamma_{1}}]$, $\ell_{\gamma_{1}},s^{*}]$, and $s^{*},\ell_{\gamma}]$, where $s^{*}$ is defined in (3.39) below.
At first, we check that
$d(\gamma(s),\partial V(t)\setminus\partial\Omega)\geq\frac{\sigma_{*}}{\sigma^{*}}s\quad\forall s\in[0,\ell_{\gamma_{1}}].$ (3.36)
To this end, fix $s\in[0,\ell_{\gamma_{1}}]$ and let $\rho:[0,\ell_{\rho}]\to\overline{\Omega}$ be any arc-length parametrized curve connecting $\gamma(s)$ to $\partial V(t)\setminus\partial\Omega$, namely, such that $\rho(0)\in\partial V(t)\setminus\partial\Omega$ and $\rho(\ell_{\rho})=\gamma(s)$. The curve resulting from following $\rho(r)$ for $r\in[0,\ell_{\rho}]$ and then $\gamma(r-\ell_{\rho}+s)$ for $r\in(\ell_{\rho},\ell_{\gamma_{1}}+\ell_{\rho}-s]$ connects $\partial V(t)$ to $\overline{V_{0}}$. Hence, formula (3.14) and the fact that $v(\cdot)=t$ on $\partial V(t)\setminus\partial\Omega$ give
$t\leq\int_{0}^{\ell_{\rho}}\sigma(\rho(r),\rho^{\prime}(r))\,{{\rm d}r}+\int_{\ell_{\rho}}^{\ell_{\gamma_{1}}+\ell_{\rho}-s}\sigma(\gamma(r-\ell_{\rho}+s),\gamma^{\prime}(r-\ell_{\rho}+s))\,{{\rm d}r}$
$\quad=\int_{0}^{\ell_{\rho}}\sigma(\rho(r),\rho^{\prime}(r))\,{{\rm d}r}+\int_{s}^{\ell_{\gamma_{1}}}\sigma(\gamma(r),\gamma^{\prime}(r))\,{{\rm d}r}.$ (3.37)
(Note that on $\partial V(t)\cap\partial\Omega$ we cannot conclude that $v(\cdot)=t$, so that the above inequality need not hold for $x\in\partial V(t)\cap\partial\Omega$). On the other hand
$\int_{0}^{s}\sigma(\gamma(r),\gamma^{\prime}(r))\,{{\rm d}r}+\int_{s}^{\ell_{\gamma_{1}}}\sigma(\gamma(r),\gamma^{\prime}(r))\,{{\rm d}r}=\int_{0}^{\ell_{\gamma_{1}}}\sigma(\gamma(r),\gamma^{\prime}(r))\,{{\rm d}r}=v(x)<t.$ (3.38)
Putting the two inequalities (3.37) and (3.38) together we obtain
$\int_{0}^{s}\sigma(\gamma(r),\gamma^{\prime}(r))\,{{\rm d}r}<\int_{0}^{\ell_{\rho}}\sigma(\rho(r),\rho^{\prime}(r))\,{{\rm d}r}.$
By using the bounds (3.10) on $\sigma$, the latter gives
$\sigma_{*}{s}<\sigma^{*}\ell_{\rho},$
implying that the length $\ell_{\rho}$ of any curve in $\overline{\Omega}$ connecting $\gamma(s)$ to $\partial V(t)\setminus\partial\Omega$ is strictly bounded from below by $\sigma_{*}s/\sigma^{*}$. Taking $\rho$ to be a straight curve in $\overline{\Omega}$ from $\gamma(s)$ to $x\in\partial V(t)\setminus\partial\Omega$ so that $|\gamma(s)-x|=d(\gamma(s),\partial V(t)\setminus\partial\Omega)$, this proves that $d(\gamma(s),\partial V(t)\setminus\partial\Omega)=\ell_{\rho}\geq\sigma_{*}s/\sigma^{*}$, giving the lower bound (3.36). Eventually, the lower bound (3.35) holds for $s\in[0,\ell_{\gamma_{1}}]$ as $\overline{\kappa}\leq\sigma_{*}/\sigma^{*}$ (recall that $\kappa_{0}\leq 1$).
Figure 2. The curve $\gamma$ is obtained by concatenating $\gamma_{1}$ and $\gamma_{2}$. The twisted cone around $\gamma_{1}$ is described by (3.36), whereas the one around $\gamma_{2}$ follows from $V_{0}$ being a John domain.
Consider now the case $s\in[\ell_{\gamma_{1}},s^{*}]$ with
$s^{*}:=\min\left\{\ell_{\gamma_{1}}\left(1+\frac{\sigma_{*}}{2\sigma^{*}}\right),\ell_{\gamma}\right\}.$ (3.39)
In particular, $s-\ell_{\gamma_{1}}\leq s^{*}-\ell_{\gamma_{1}}\leq\sigma_{*}\ell_{\gamma_{1}}/(2\sigma^{*})$. This implies that
$|y-\gamma(s)|$ $=|y-\gamma_{2}(s-\ell_{\gamma_{1}})|=|\gamma_{2}(0)-\gamma_{2}(s-\ell_{\gamma_{1}})|\leq s-\ell_{\gamma_{1}}\leq{\sigma_{*}\ell_{\gamma_{1}}}/{(2\sigma^{*})},$
where we also used that $|\gamma_{2}(0)-\gamma_{2}(s-\ell_{\gamma_{1}})|\leq s-\ell_{\gamma_{1}}$, as $\gamma_{2}$ is arc-length parametrized. As $d(y,\partial V(t)\setminus\partial\Omega)\geq{\sigma_{*}\ell_{\gamma_{1}}}/{\sigma^{*}}$ from (3.36), the triangle inequality allows us to conclude that
$\frac{\sigma_{*}\ell_{\gamma_{1}}}{\sigma^{*}}\leq d(y,\partial V\setminus\partial\Omega)\leq d(\gamma(s),\partial V\setminus\partial\Omega)+|y-\gamma(s)|\leq d(\gamma(s),\partial V\setminus\partial\Omega)+\frac{\sigma_{*}\ell_{\gamma_{1}}}{2\sigma^{*}}.$
This implies that $d(\gamma(s),\partial V(t)\setminus\partial\Omega))\geq{\sigma_{*}\ell_{\gamma_{1}}}/{(2\sigma^{*})}$ and (3.35) holds for $s\in[\ell_{\gamma_{1}},s^{*}]$, as well as
[…]
We are hence left with considering the case $s\in[s^{*},\ell_{\gamma}]$. Note that this case is only relevant if $s^{*}<\ell_{\gamma}$, hence for $s^{*}=\ell_{\gamma_{1}}(1+\sigma_{*}/(2\sigma^{*}))$. As $s>\ell_{\gamma_{1}}$, we have that $\gamma(s)\in V_{0}$. We use that $\overline{V_{0}}$ is a John domain with John center $x_{0}$ and John constant $\kappa_{0}$ to obtain
$d(\gamma(s),\partial V(t)\setminus\partial\Omega)\geq d(\gamma(s),\partial V_{0})=d(\gamma_{2}(s-\ell_{\gamma_{1}}),\partial V_{0})\geq\kappa_{0}(s-\ell_{\gamma_{1}}).$
The lower bound (3.35) for $s\in[s^{*},\ell_{\gamma}]$ follows then from
[…]
Step 2: The curve $\widetilde{\gamma}$.
To prove that $V(t)$ is a John domain, we need to complement (3.35) by providing a lower bound for the distance of $\gamma(s)$ from $\partial V(t)\cap\partial\Omega$. This however may require to modify the curve $\gamma$, as a priori $\gamma$ may get close and even touch $\partial V(t)$ at points in $\partial V(t)\cap\partial\Omega$, see Figure 3. We hence define a new curve $\widetilde{\gamma}$ by displacing $\gamma$ in the inward normal direction in a neighborhood of the boundary $\partial V(t)\cap\partial\Omega$. This may also require to reduce the John constant.
Figure 3. Construction of the curve $\widetilde{\gamma}$ (Step 2). The curve $\gamma$ touches the boundary $\partial\Omega$. Within the $\varepsilon_{0}$-neighborhood of $\partial\Omega$ (indicated by the dashed curve), the curve $\widetilde{\gamma}$ is obtained by displacing $\gamma$ in the inward normal direction. This allows to fit a twisted cone around $\widetilde{\gamma}$ contained in $V(t)$.
Let us start by recalling that the projection $\pi:(\partial\Omega+B_{\overline{\varepsilon}})\cap\overline{\Omega}\to\partial\Omega$ and the normal outward-pointing vector field $\nu:\partial\Omega\to{\mathbb{S}}^{n-1}$ are well-defined and Lipschitz continuous [39, Cor. 2] and set $\eta:=-\nu\circ\pi$, which is Lipschitz continuous, as well.
Fix $0<\varepsilon_{0}<\min\{\overline{\varepsilon},d(x_{0},\partial\Omega)\}$ and set $\widetilde{\kappa}>0$ so small that
$\widetilde{\kappa}<\min\left\{\frac{\overline{\kappa}}{2},\frac{\varepsilon_{0}}{T},\frac{1}{2(\|\nabla\eta\|_{\infty}T+T/\varepsilon_{0}+1)}\right\}.$ (3.40)
We define $\widetilde{\gamma}:[0,\ell_{\gamma}]\to\Omega$ as
$\widetilde{\gamma}(s):=\gamma(s)+\eta(\gamma(s))\big(\varepsilon_{0}-d(\gamma(s),\partial\Omega)\big)^{+}\frac{\widetilde{\kappa}}{\varepsilon_{0}}s.$
Note that $\widetilde{\gamma}$ is Lipschitz continuous. In fact, for all $s,\,t\in(0,\ell_{\gamma})$ we have that
$\widetilde{\gamma}(t)-\widetilde{\gamma}(s)=\gamma(t)-\gamma(s)$
$\quad+\big(\eta(\gamma(t))-\eta(\gamma(s))\big)\big(\varepsilon_{0}-d(\gamma(t),\partial\Omega)\big)^{+}\frac{\widetilde{\kappa}}{\varepsilon_{0}}t$
$\quad+\eta(\gamma(s))\Big(\big(\varepsilon_{0}-d(\gamma(t),\partial\Omega)\big)^{+}-\big(\varepsilon_{0}-d(\gamma(s),\partial\Omega)\big)^{+}\Big)\frac{\widetilde{\kappa}}{\varepsilon_{0}}t$
$\quad+\eta(\gamma(s))\big(\varepsilon_{0}-d(\gamma(s),\partial\Omega)\big)^{+}\frac{\widetilde{\kappa}}{\varepsilon_{0}}(t-s).$
This ensures that
$|\widetilde{\gamma}(t)-\widetilde{\gamma}(s)|\leq|\gamma(t)-\gamma(s)|+\widetilde{\kappa}\left(\|\nabla\eta\|_{L^{\infty}}T+T/\varepsilon_{0}\right)|\gamma(t)-\gamma(s)|+\widetilde{\kappa}|t-s|$
and, in particular,
$\|\widetilde{\gamma}^{\prime}\|_{L^{\infty}}\leq 1+\widetilde{\kappa}\big(\|\nabla\eta\|_{L^{\infty}}T+T/\varepsilon_{0}+1\big)=:M<3/2\quad\text{a.e. in $(0,\ell_{\gamma})$}.$ (3.41)
A similar argument entails that
$\|\widetilde{\gamma}^{\prime}\|_{L^{\infty}}\geq 1-\widetilde{\kappa}\big(\|\nabla\eta\|_{L^{\infty}}T+T/\varepsilon_{0}+1\big)>1/2\quad\text{a.e. in $(0,\ell_{\gamma})$}.$ (3.42)
Moreover, $\widetilde{\gamma}(0)=\gamma(0)=x$, and, using also $\gamma(\ell_{\gamma})=x_{0}$ and the fact that $\varepsilon_{0}<d(x_{0},\partial\Omega)$, we have that
$\widetilde{\gamma}(\ell_{\gamma})=x_{0}+\eta(x_{0})\big(\varepsilon_{0}-d(x_{0},\partial\Omega)\big)^{+}\frac{\widetilde{\kappa}}{\varepsilon_{0}}\ell_{\gamma}=x_{0}.$
The lower bound (3.35) and the triangle inequality allow us to control
[…]
$\quad=d(\widetilde{\gamma}(s),\partial V(t)\setminus\partial\Omega)+\big(\varepsilon_{0}-d(\gamma(s),\partial\Omega)\big)^{+}\frac{\widetilde{\kappa}}{\varepsilon_{0}}s\leq d(\widetilde{\gamma}(s),\partial V(t)\setminus\partial\Omega)+\widetilde{\kappa}s$
entailing that
$d(\tilde{\gamma}(s),\partial V(t)\setminus\partial\Omega)\geq(\overline{\kappa}-\widetilde{\kappa})s>(2\widetilde{\kappa}-\widetilde{\kappa})s=\widetilde{\kappa}s\quad\forall s\in[0,\ell_{\gamma}].$ (3.43)
Fix now $s\in[0,\ell_{\gamma}]$. If $d(\gamma(s),\partial\Omega)\geq\varepsilon_{0}$ we use the fact that $\widetilde{\kappa}<\varepsilon_{0}/T$ from (3.40) to find $d(\widetilde{\gamma}(s),\partial\Omega)\geq\varepsilon_{0}\geq\widetilde{\kappa}T\geq\widetilde{\kappa}s$. On the contrary, if $d(\gamma(s),\partial\Omega)<\varepsilon_{0}$ we have
$d(\widetilde{\gamma}(s),\partial\Omega)=d(\gamma(s),\partial\Omega)+\big(\varepsilon_{0}-d(\gamma(s),\tilde{\Omega})\big)\frac{\widetilde{\kappa}}{\varepsilon_{0}}s$
$\quad=\left(1-\frac{\widetilde{\kappa}}{\varepsilon_{0}}s\right)d(\gamma(s),\partial\Omega)+\widetilde{\kappa}s\geq\left(1-\frac{\widetilde{\kappa}}{\varepsilon_{0}}T\right)d(\gamma(s),\partial\Omega)+\widetilde{\kappa}s\stackrel{{\scriptstyle\widetilde{\kappa}<\varepsilon_{0}/T}}{{\geq}}\widetilde{\kappa}s.$
In both cases, we hence have that
$d(\widetilde{\gamma}(s),\partial\Omega)\geq\widetilde{\kappa}s\quad\forall s\in[0,\ell_{\gamma}].$ (3.44)
Putting together (3.43) and (3.44) we conclude that
$d(\widetilde{\gamma}(s),\partial V(t))=\min\big\{d(\widetilde{\gamma}(s),\partial V\setminus\partial\Omega),d(\widetilde{\gamma}(s),\partial V(t)\cap\partial\Omega)\big\}\geq\widetilde{\kappa}s\quad\forall s\in[0,\ell_{\gamma}].$ (3.45)
Step 3: Arc-length reparametrization of $\widetilde{\gamma}$.
The last step of the proof consists in reparametrizing $\widetilde{\gamma}$ by arc length: One defines $\widehat{\gamma}:[0,\ell_{\widehat{\gamma}}]\to\overline{\Omega}$ as $\widehat{\gamma}(\tau)=\widetilde{\gamma}(s(\tau))$ where $s:[0,\ell_{\widehat{\gamma}}]\to[0,\ell_{\gamma}]$ is the inverse function of $s\mapsto\tau(s):=\int_{0}^{s}|\widetilde{\gamma}^{\prime}(r)|\,{\rm d}r$. Note that $s$ is Lipschitz continuous due to (3.42). Using (3.41), for all $\tau\in[0,\ell_{\widehat{\gamma}}$] we have that
[…]
Hence, setting $\kappa:=\widetilde{\kappa}/M$ we obtain
[…]
This implies that $V(t)$ is a John domain with John center $x_{0}$ and John constant
$\kappa(T)=\frac{\widetilde{\kappa}}{1+\widetilde{\kappa}(\|\nabla\eta\|_{L^{\infty}}T+T/\varepsilon_{0}+1)}>0.$
In particular, $\kappa(T)$ depends on $\kappa_{0}$, $\sigma_{*}$, $\sigma^{*}$, $\varepsilon_{0}$, $\|\nabla\eta\|_{L^{\infty}}$, and $T$, and $\kappa(T)\to 0$ as $T\to\infty$. Moreover, $\varepsilon_{0}$ depends on $\overline{\varepsilon}$ and $d(x_{0},\partial\Omega)$.
Step 4: Lipschitz continuity of $t\mapsto V(t)$ with respect to the Hausdorff distance.
Assume that $0<s<t\leq T$, choose $x\in V(t)\setminus\overline{V(s)}$, and let $\gamma_{x}\in{\rm Lip}_{x}$ be optimal for $x$ with $\gamma_{x}(\cdot)\in\overline{\Omega}$. For $\varepsilon>0$ small, fix $\tau\in(0,1)$ with $0<s-\varepsilon<v(\gamma_{x}(\tau))<s$, so that $\gamma_{x}(\tau)\in V(s)$. Such $\tau$ exists, as $s\in[0,1]\mapsto v(\gamma_{x}(s))$ is continuous, $v(\gamma_{x}(0))=0$, and $v(\gamma_{x}(1))=v(x)>s>0$. We have that
$\sigma_{*}d(x,V(s))\leq\sigma_{*}|x-\gamma_{x}(\tau)|\leq\sigma_{*}d_{\overline{\Omega}}(x,\gamma_{x}(\tau))\leq\sigma_{*}\int_{\tau}^{1}|\gamma_{x}^{\prime}(r)|\,{\rm d}r$
[…]
$\quad\leq v(x)-v(\gamma_{x}(\tau))<t-s+\varepsilon.$
By taking the infimum for $\varepsilon>0$ and then the supremum on $x\in V(t)$ we obtain (3.32).
Step 5: Hölder continuity of $t\mapsto V(t)$ with respect to the $L^{1}$ distance.
Let $0<s<t\leq T$ and recall that $V(t)$ is a John domain with respect to $x_{0}$ with John constant $\kappa(T)$. Owing to [43, Cor. 6.4], see also [49, Cor. 2], the Minkowski (or box-counting) dimension of $\partial V(t)$ is smaller than $n$. More precisely, the minimal number $N(r)$ of balls with radius $r>0$ needed to cover $\partial V(t)$ can be bounded as
$N(r)\leq cr^{\mu-n}$ (3.47)
where $\mu\in(0,1)$ and $c>0$ depend on
$\ell_{\rm max}(T):=\sup_{x\in\overline{V(T)}}d_{\overline{\Omega}}(x,x_{0})\leq\frac{d(x_{0},\partial V_{0})}{\kappa_{0}}+\frac{T}{\sigma_{*}}$ (3.48)
and $\kappa(T)$ only.
Fix $r:=2(t-s)/\sigma_{*}$, and let $K$ be a finite covering of $\partial V(t)$ with $N(r)$ balls of radius $r$. The bound (3.32) guarantees that $d_{H}(V(t),V(s))\leq(t-s)/\sigma_{*}=r/2$ so that $V(t)\setminus V(s)\subset K$. Using (3.47) we conclude that
$|V(t)\setminus V(s)|\leq|K|\leq N(r)|B_{1}|r^{n}\leq c|B_{1}|\left(\frac{2(t-s)}{\sigma_{*}}\right)^{\mu}$
where $\mu\in(0,1)$ and $c$ depend on $\ell_{\rm max}(T)$ and the John constant $\kappa(T)$ of $V(t)$ only. We obtain (3.33) by letting $\mu(T)=\mu$ and $c(T)=c|B_{1}|2^{\mu}\sigma_{*}^{-\mu}$. ∎
4. Free boundary problem: Proof of Theorem 2.1
We start by recalling a uniform Poincaré estimate from [8, Lemma 3.1] and [8, Theorem 5.1].
Proposition 4.1 (Uniform Poincaré inequality).
Assume (2.1)–(2.4) and define
[…]
$\qquad\qquad\text{with John constant $\kappa\leq\kappa_{0}$}\}.$ (4.1)
Then, there exists a constant $c_{\mathcal{V}}>0$ depending on $\kappa_{0}$, $d(x_{0},\partial V_{0})$, and $R$ such that
$\|u\|_{H^{1}_{\Gamma}(V)}\leq c_{\mathcal{V}}\|\nabla u\|_{L^{2}(V;{\mathbb{R}}^{n})}\quad\forall V\in\mathcal{V},\ \forall u\in H^{1}_{\Gamma}(V).$ (4.2)
Proof of Theorem 2.1.
We use an iteration argument, which is divided into steps. We first consider the elliptic problem (2.12) for given $v$ (Step 1). Then, by solving (2.13) for a given $u$ we define the iteration (Step 2) and discuss its convergence (Step 3). Specifically, we prove that one can pass to the iteration limit along subsequences both in the elliptic problem (Step 4) and in the Hamilton–Jacobi problem (Step 5). Eventually, we prove that the stronger (2.14) implies that $v$ is a constrained viscosity solution (Step 6).
Step 1: Solution to the elliptic problem, given the growth. Assume to be given $v\in C(\overline{\Omega})$ such that for all $T>0$ and all $t\in[0,T]$ the level set $V_{0}\subset V(t)=\{x\in\overline{\Omega}\ |\ v(x)<t\}$ is a John domain with respect to $x_{0}\in V_{0}$ with John constant $\kappa(T)$. Owing to the standard Lax–Milgram Lemma, for all fixed $t>0$ one can find a unique $u(\cdot,t)\in H^{1}_{\Gamma}(V(t))$ solving
$\int_{V(t)}\nabla u(x,t)\cdot\nabla w(x)\,{\rm d}x=\int_{V(t)}w(x)\,{\rm d}x\quad\forall w\in H^{1}_{\Gamma}(V(t)).$ (4.3)
Moroever, one can prove that the ensuing $u:Q_{v}=\cup_{t>0}V(t)\times\{t\}\to{\mathbb{R}}$ is measurable. By choosing $w=u(\cdot,t)$ in (4.3) and using the uniform Poincaré inequality (4.2) one gets
$\|u(\cdot,t)\|_{H^{1}_{\Gamma}(V(t))}\leq c_{2}(T)\quad\forall t\in[0,T],\ \forall T>0$ (4.4)
where the increasing function $c_{2}:{\mathbb{R}}_{+}\to{\mathbb{R}}_{+}$ depends on $T$ via $\kappa(T)$.
Step 2: Construction of the iteration. Under assumptions (2.5)–(2.7) and (2.9), the Hamiltonian $(x,p)\in\overline{\Omega}\times{\mathbb{R}}^{n}\mapsto H(x,0,p)$ fulfills (3.5)–(3.7) and (3.9). One implication in (3.8) is directly implied by (2.8), while the converse implication follows by continuity. As (2.1)–(2.2) are assumed to hold, owing to Proposition 3.3 we find $v_{0}\in C(\overline{\Omega})$ solving
$H(x,0,\nabla v_{0})=0\quad\text{in}\ \Omega\setminus\overline{V_{0}},$
$v_{0}=0\quad\text{in}\ V_{0}$
given by (3.11), with $\sigma$ defined from $H(x,0,\cdot)$. Proposition 3.8 ensures that, for all $T>0$ and $t\in[0,T]$, the level sets $\{x\in\overline{\Omega}\ |\ v_{0}(x)<t\}$ are John domains with respect to $x_{0}$ with John constant $\kappa(T)$. We can hence argue as in Step 1 and find the unique solution $u_{0}$ to (4.3) with $V(t)$ replaced by $\{x\in\overline{\Omega}\ |\ v_{0}(x)<t\}$.
Assume now to be given $v_{j}\in C(\overline{\Omega})$ and $Ku_{j}\in C({\mathbb{R}}^{n}\times{\mathbb{R}}_{+})$. As the composition $x\in\overline{\Omega}\mapsto Ku_{j}(x,v_{j}(x))$ is continuous, the Hamiltonian $(x,p):\overline{\Omega}\times{\mathbb{R}}^{n}\mapsto H(x,Ku_{j}(x,v_{j}(x)),p)$ fulfills (3.5)–(3.9). Again (3.5)–(3.7), (3.9), and one implication in (3.8) follow directly from (2.5)–(2.9) and continuity implies the converse implication in (3.8). Proposition 3.3 guarantees that a viscosity solution $v_{j+1}\in C(\overline{\Omega})$ to
$H(x,Ku_{j}(x,v_{j}(x)),\nabla v_{j+1})=0\quad\text{in}\ \Omega\setminus\overline{V_{0}},$ (4.5)
$v_{j+1}=0\quad\text{in}\ V_{0}.$ (4.6)
is given by the representation formula
$v_{j+1}(x)=\min\Bigg\{\int_{0}^{1}\sigma\big(\gamma,Ku_{j}(\gamma,v_{j}(\gamma)),\gamma^{\prime}\big)\ {{\rm d}s}\ \Big|\$$\gamma\in{\rm Lip}_{x},\ \gamma(\cdot)\in\overline{\Omega}\Bigg\}$ (4.7)
for all $x\in\overline{\Omega}$, where we have used the support function
$\sigma(x,u,q):=\sup\{q\cdot p\ |\ H(x,u,p)\leq 0\}\quad\forall(x,u,q)\in\overline{\Omega}\times{\mathbb{R}}^{n}\times{\mathbb{R}}^{n}.$
Under assumptions (2.5)–(2.7) and (2.9), the statement of Lemma 3.1 can be readily extended to cover the case of $\sigma$ depending on $u$, as well. In particular, we have that
$\sigma\ \ \text{is continuous and}\ \ \sigma(x,u,\cdot)\ \ \text{is convex \ $\forall(x,u)\in\overline{\Omega}\times{\mathbb{R}}$}.$ (4.8)
Again by Proposition 3.8, all sublevels $V_{j+1}(t)=\{x\in\overline{\Omega}:v_{j+1}(x)<t\}$ are John domains with respect to $x_{0}$ with John constant $\kappa(T)$ for $t\in[0,T]$. In fact, all Hamiltonians $(x,p)\mapsto H(x,Ku_{j}(x,v_{j}(x)),p)$ fulfill (3.5)–(3.9) with the same $0<\sigma_{*}\leq\sigma^{*}$ and $\kappa(T)$ just depends on $\sigma_{*}$, $\sigma^{*}$, and $T$, in addition to $\Omega$ and $V_{0}$ (via $\overline{\varepsilon}$, $\|\nabla\eta\|_{L^{\infty}}$, $\kappa_{0}$, and $d(x_{0},\partial\Omega)$). We can hence uniquely solve (4.3) with $V(t)$ replaced by $V_{j+1}(t)$ to obtain $u_{j+1}$ fulfilling (4.4). In particular, one has that $Ku_{j+1}\in C({\mathbb{R}}^{n}\times{\mathbb{R}}_{+})$.
By iterating this construction we obtain sequences $(v_{j})_{j}\in C(\overline{\Omega})$ and $(u_{j})_{j}$ for $j\geq 2$ solving
$\int_{V_{j}(t)}\nabla u_{j}(x,t)\cdot\nabla w(x)\,{\rm d}x=\int_{V_{j}(t)}w(x)\,{\rm d}x\quad\forall w\in H^{1}_{\Gamma}(V_{j}(t)),\ \text{for a.e.}\ t>0,$ (4.9)
$v_{j}(x)=\min\Bigg\{\int_{0}^{1}\sigma\big(\gamma,Ku_{j-1}(\gamma,v_{j-1}(\gamma)),\gamma^{\prime}\big)\ {{\rm d}s}:\gamma\in{\rm Lip}_{x},\gamma(\cdot)\in\overline{\Omega}\Bigg\}\quad\forall x\in\overline{\Omega},$ (4.10)
and fulfilling the estimates, see (3.30)–(3.31) and (4.4),
$\|u_{j}(\cdot,t)\|_{H^{1}_{\Gamma}(V(t))}\leq c_{2}(T)\quad\forall t\in[0,T],\ \forall T>0,$ (4.11)
$|\nabla v_{j}|\leq\sigma^{*}L\quad\text{a.e. in}\ \Omega,$ (4.12)
$v_{j}(x)\leq c_{1}(R)\quad\forall x\in B_{R},\ \forall R>0,\ \forall j\geq 2.$ (4.13)
Moreover, for all $x\in V_{j}(t)$ and $y\in\overline{V_{0}}$ such that $|x-y|=d(x,\partial V_{0})$, letting $\gamma:[0,\ell_{\gamma}]\to\overline{V_{0}}$ be arc-parametrized with $d(\gamma(s),\partial V_{0})\geq\kappa_{0}s$ for all $s\in[0,\ell_{\gamma}]$, we get
$|x|\leq|x-y|+|y-x_{0}|+|x_{0}|\leq\frac{t}{\sigma_{*}}+\ell_{\gamma}+|x_{0}|\leq\frac{t}{\sigma_{*}}+\frac{d(x_{0},\partial V_{0})}{\kappa_{0}}+|x_{0}|.$
This proves that
$V_{j}(t)\subset B_{c_{3}(t)}\quad\forall t>0$ (4.14)
with $c_{3}(t):=t/\sigma_{*}+d(x_{0},\partial V_{0})/\kappa_{0}+|x_{0}|$.
To proceed, we first first recall (3.32)–(3.33) from Proposition 3.8, which are uniform with respect to $v_{j}$, namely,
$d_{H}(V_{j}(t_{1}),V_{j}(t_{2}))\leq\frac{1}{\sigma_{*}}|t_{1}-t_{2}|\quad\forall t_{1},\,t_{2}>0,\ \forall j\geq 2,$ (4.15)
$|V_{j}(t_{2})\setminus V_{j}(t_{1})|\leq c_{H}(T)|t_{2}-t_{1}|^{\mu(T)}\quad\forall 0<t_{1}<t_{2}<T.$ (4.16)
Step 3: Convergence. Set $\Omega_{R}:=\Omega\cap B_{R}$ for all $R>0$. Denote by $\overline{u_{j}}$ and $\overline{\nabla u_{j}}$ the trivial extensions to the whole ${\mathbb{R}}^{n}$ of $u_{j}$ and $\nabla u_{j}$, respectively. Owing to estimates (4.11)–(4.12), by a diagonal extraction argument we find not relabeled subsequences such that
$\overline{u_{j}}\to\overline{u}\quad\text{weakly* in}\ L^{\infty}(0,T;L^{2}({\mathbb{R}}^{n}))\quad\forall T>0,$ (4.17)
$\overline{\nabla u_{j}}\to G\quad\text{weakly* in}\ L^{\infty}(0,T;L^{2}({\mathbb{R}}^{n};{\mathbb{R}}^{n}))\quad\forall T>0,$ (4.18)
$v_{j}\to v\quad\text{strongly in}\ C(\overline{\Omega_{R}})\ \text{and}\ \text{weakly* in}\ W^{1,\infty}(\Omega_{R})\quad\forall R>0.$ (4.19)
Set $u$ to be the restriction to $Q_{v}$ of $\overline{u}$.
Fix $t>0$ and consider the two level sets $V(t)$ and $V_{j}(t)$. Choose $R>c_{3}(t)$, so that $V_{j}(t)\subset\Omega_{R}$ from (4.14). Assume that $V(t)\setminus V_{j}(t)$ is not empty and fix $x\in V(t)\setminus V_{j}(t)$. As $v_{j}\to v$ uniformly on $\overline{\Omega}\cap B_{R}$, setting $\delta=\|v-v_{j}\|_{L^{\infty}(\overline{\Omega_{R}})}$ one readily proves that $V_{j}(t-\delta)\subset V(t)\subset V_{j}(t+\delta)$. Indeed, given $x\in V_{j}(t-\delta)$ and $y\in VV(t)$ one has $v(x)\leq v_{j}(x)+\delta<t+\delta$ and $v_{n}(y)\leq v(y)+\delta<t+\delta$. In particular, this guarantees that
$d_{H}(V(t),V_{j}(t))$ $=\max\left\{\sup_{x\in V(t)}\inf_{y\in V_{j}(t)}|x-y|,\sup_{x\in V_{j}(t)}\inf_{y\in V(t)}|x-y|\right\}$
$\leq\max\left\{\sup_{x\in V_{j}(t+\delta)}\inf_{y\in V_{j}(t)}|x-y|,\sup_{x\in V_{j}(t)}\inf_{y\in V_{j}(t-\delta)}|x-y|\right\}$
[…]
In other words, owing to convergence (4.19) we have that
$d_{H}(V_{j}(t),V(t))\to 0\quad\text{locally uniformly w.r.t.}\ t>0.$ (4.21)
This ensures the convergence of the corresponding characteristic functions: Take any $(x,t)\in\overline{\Omega}\times{\mathbb{R}}_{+}$ with $v(x)<t$ ($v(x)>t$, respectively). One has that $v_{j}(x)<t$ ($v_{j}(x)>t$) for $j$ large enough. Hence, $(x,t)\in\overline{\Omega}\times{\mathbb{R}}_{+}\to I_{j}(x,t):=1_{V_{j}(t)}(x)$ (characteristic functions) converges pointwise to $(x,t)\in\overline{\Omega}\times{\mathbb{R}}_{+}\to I(x,t)=1_{V(t)}(x)$ at all points $(x,t)$ with $v(x)\not=t$. On the other hand, the set $\{(x,t)\in\overline{\Omega}\times{\mathbb{R}}_{+}\>|\>v(x)=t\}$ is negligible, so that the above convergence holds almost everywhere. We conclude that
$I_{j}\to I\quad\text{strongly in}\ L^{p}_{\rm loc}(\Omega\times{\mathbb{R}}_{+})\quad\forall p\in[1,\infty).$ (4.22)
Fix now an arbitrary $(x,t)\in Q_{v}$ and let $\delta>0$ be small, in such a way that $A:=[(t-\delta,t+\delta)\times B_{\delta}(x)\subset\subset Q_{v}$. Owing to convergence (4.19) one has that $A\subset Q_{v_{j}}$ for $j$ large enough. By restricting to $A$ we hence have that $u_{j}\to u$ weakly star in $L^{\infty}(t-\delta,t+\delta;H^{1}(B_{\delta}(x))$, which entails that $G=\nabla u$ in $A$. By exhausting $Q_{v}$, this proves that
$G=\nabla u\quad\text{a.e. in}\ Q_{v}.$ (4.23)
Step 4: Passage to the limit in the elliptic problem. Let $D\subset H^{1}_{\Gamma}(\Omega)$ be dense and countable. For all $w\in D$, let $J_{w}\subset{\mathbb{R}}_{+}$ be such that ${\mathbb{R}}_{+}\setminus J_{w}$ are Lebesgue points of
$t\in{\mathbb{R}}_{+}\mapsto\int_{V(t)}\nabla u(x,t)\cdot\nabla w(x)\,{\rm d}x=\int_{\Omega}\nabla u(x,t)\cdot\nabla w(x)\,I(x,t)\,{\rm d}x.$ (4.24)
Set $t\in{\mathbb{R}}_{+}\setminus J$ for $J:=\cup_{w\in D}J_{w}$, choose $w\in D$ in (4.9), and integrate over $(t-\delta,t)$ for some small $\delta\in(0,t)$. For $R>c_{3}(t)$ one has
$\frac{1}{\delta}\int_{t-\delta}^{t}\!\int_{\Omega_{R}}\overline{\nabla u_{j}}(x,s)\cdot\nabla w(x)\,I_{j}(x,s)\,{\rm d}x\,{\rm d}s=\frac{1}{\delta}\int_{t-\delta}^{t}\!\int_{V_{j}(s)}{\nabla u_{j}}(x,s)\cdot\nabla w(x)\,{\rm d}x\,{\rm d}s$
$\quad=\frac{1}{\delta}\int_{t-\delta}^{t}\!\int_{V_{j}(s)}w(x)\,{\rm d}x\,{\rm d}s=\frac{1}{\delta}\int_{t-\delta}^{t}\!\int_{\Omega_{R}}w(x)\,I_{j}(x,s)\,{\rm d}x\,{\rm d}s.$
Convergences (4.18) and (4.22) and the identification (4.23) allow to pass to the limit and get
$\frac{1}{\delta}\int_{t-\delta}^{t}\!\int_{V(s)}\nabla u(x,s)\cdot\nabla w(x)\,{\rm d}x\,{\rm d}s=\frac{1}{\delta}\int_{t-\delta}^{t}\!\int_{\Omega_{R}}\nabla u(x,s)\cdot\nabla w(x)\,I(x,s)\,{\rm d}x\,{\rm d}s$
$\quad=\frac{1}{\delta}\int_{t-\delta}^{t}\!\int_{\Omega_{R}}w(x)\,I(x,s)\,{\rm d}x\,{\rm d}s=\frac{1}{\delta}\int_{t-\delta}^{t}\!\int_{V(s)}w(x)\,{\rm d}x\,{\rm d}s.$ (4.25)
In order to pass to the limit as $\delta\to 0$, one can handle the above right-hand side as follows
$\left|\frac{1}{\delta}\int_{t-\delta}^{t}\!\int_{V(s)}w(x)\,{\rm d}x\,{\rm d}s-\int_{V(t)}w(x)\,{\rm d}x\right|=\left|\frac{1}{\delta}\int_{t-\delta}^{t}\!\int_{V(t)\setminus V(s)}w(x)\,{\rm d}x\,{\rm d}s\right|$
$\quad\leq\frac{1}{\delta}\int_{t-\delta}^{t}\!\int_{V(t)\setminus V(s)}|w(x)|\,{\rm d}x\,{\rm d}s\leq|V(t)\setminus V(t-\delta)|^{1/2}\|w\|_{L^{2}(\Omega)}$
[…]
Using the fact that $t$ is a Lebesgue point, see (4.24), and passing to the limit as $\delta\to 0$ in (4.25) we hence get
$\int_{V(t)}\nabla u(x,t)\cdot\nabla w(x)\,{\rm d}x=\int_{V(t)}w(x)\,{\rm d}x.$
Relation (2.12) follows from the density of $D$ in $H^{1}(\Omega)$ by recalling that $|J|=0$.
Step 5: Passage to the limit in the Hamilton–Jacobi problem. Owing to convergence (4.17), by applying the mollification we have that
$Ku_{j}\to Ku\quad\text{locally uniformly in}\ \overline{\Omega}\times{\mathbb{R}}_{+}.$ (4.26)
In addition, using the fact that
$\partial_{t}Ku_{j}(x,t)=k(0)\int_{{\mathbb{R}}^{n}}\phi(x-y)\overline{u_{j}}(y,t)\,{\rm d}y+\int_{0}^{t}\!\!\int_{{\mathbb{R}}^{n}}k^{\prime}(t-s)\phi(x-y)\overline{u_{j}}(y,s)\,{\rm d}y\,{\rm d}s$
we can prove that $Ku_{j}$ are locally Lipschitz continuous in time, namely,
$\|Ku_{j}(\cdot,t_{1})-Ku_{j}(\cdot,t_{2})\|_{L^{\infty}(\overline{\Omega})}\leq\|\partial_{t}Ku_{j}\|_{L^{\infty}(\overline{\Omega}\times(t_{1},t_{2}))}|t_{1}-t_{2}|$
$\quad\leq\big(|k(0)|+\|k^{\prime}\|_{L^{1}(t_{1},t_{2})}\big)\|\phi\|_{L^{2}({\mathbb{R}}^{n})}\|\overline{u}_{j}\|_{L^{\infty}(t_{2},t_{1};L^{2}({\mathbb{R}}^{n}))}|t_{1}-t_{2}|$
$\quad\leq\big(|k(0)|+\|k^{\prime}\|_{L^{1}(t_{1},t_{2})}\big)\|\phi\|_{L^{2}({\mathbb{R}}^{n})}c_{2}(T)|t_{1}-t_{2}|$ (4.27)
for all $0<t_{1}<t_{2}<T$. This allows to control
$\sup_{x\in\overline{\Omega_{R}}}|Ku_{j}(x,v_{j}(x))-Ku(x,v(x))|$
$\quad\leq\sup_{x\in\overline{\Omega_{R}}}\big(|Ku_{j}(x,v_{j}(x))-Ku_{j}(x,v(x))|+|Ku_{j}(x,v(x))-Ku(x,v(x))|\big)$
$\quad\leq\big(|k(0)|+\|k^{\prime}\|_{L^{1}(t_{1},t_{2})}\big)\|\phi\|_{L^{2}({\mathbb{R}}^{n})}c_{2}(c_{1}(R))\,\|v-v_{j}\|_{L^{\infty}(\overline{\Omega_{R}})}$
$\qquad+\|Ku_{j}-Ku\|_{L^{\infty}(\overline{\Omega_{R}}\times[0,c_{1}(R)])}$ (4.28)
for all $R>0$. Convergences (4.19) and (4.26) entail that $Ku_{j}(\cdot,v_{j}(\cdot))\to Ku(\cdot,v(\cdot))$ locally uniformly. By standard arguments (see, e.g., Step 4 in the proof of Proposition 3.3) this implies that $v$ is a viscosity solution to (1.1). As $v_{j}=0$ on $V_{0}$, condition (1.3) follows.
In order to check the validity of the representation formula (2.13), set $x\in\overline{\Omega}$. For any $\gamma\in{\rm Lip}_{x}$ with $\gamma(\cdot)\in\overline{\Omega}$, using (4.28) we get that
$Ku_{j-1}(\gamma,v_{j-1}(\gamma))\to Ku(\gamma,v(\gamma))\quad\text{uniformly in}\ [0,1].$ (4.29)
Using the continuity of $\sigma$ from (4.8) and dominated convergence we hence have that
$v(x)=\lim_{j\to\infty}v_{j}(x)\leq\lim_{j\to\infty}\int_{0}^{1}\sigma(\gamma,Ku_{j-1}(\gamma,v_{j-1}(\gamma)),\gamma^{\prime})\,{\rm d}s$
$\quad=\int_{0}^{1}\sigma(\gamma,Ku(\gamma,v(\gamma)),\gamma^{\prime})\,{\rm d}s.$ (4.30)
To conclude for (2.13), we need to find an optimal curve realizing the equality in (4.30). We argue as in the proof of Proposition 3.3. Let $\gamma_{xj}\in{\rm Lip}_{x}$ with $\gamma_{xj}(\cdot)\in\overline{\Omega}$ be optimal curves for $x$ under $\sigma(\cdot,Ku_{j-1}(\cdot,v_{j-1}(\cdot)),\cdot)$. Define $s_{j}:[0,1]\to[0,1+v_{j}(x)]$ to be $s_{j}(t):=t+\int_{0}^{t}|\gamma_{xj}^{\prime}(r)|\,{\rm d}r$ for $t\in[0,1]$, denote by $t_{j}:[0,1+v_{j}(x)]\to[0,1]$ its inverse, and define $\widehat{\gamma}_{xj}(r):=\gamma_{xj}(t_{j}((1+v_{j}(x))r))$ for $r\in[0,1]$. Then, $\widehat{\gamma}_{xj}\in{\rm Lip}_{x}$ is optimal and $|\widehat{\gamma}_{xj}^{\prime}|\leq 1+v_{j}(x)$ a.e. Moreover, $\widehat{\gamma}_{xj}(1)=x$ and $v_{j}(x)\to v(x)$, so that the sequence $(\widehat{\gamma}_{xj})_{j}$ is uniformly bounded in $W^{1,\infty}(0,1;{\mathbb{R}}^{n})$. By extracting without relabeling we have that $\widehat{\gamma}_{xj}\to\gamma$ uniformly in $[0,1]$ and weakly* in $W^{1,\infty}(0,1;{\mathbb{R}}^{n})$. In particular, $\gamma(1)=\lim_{j\to\infty}\widehat{\gamma}_{xj}(1)=x$ and $\gamma(0)\in\overline{V_{0}}$ as $\overline{V_{0}}\ni\widehat{\gamma}_{xj}(0)\to\gamma(0)$ and $\overline{V_{0}}$ is closed, hence $\gamma\in{\rm Lip}_{x}$. We hence conclude that
$v(x)$ $=\lim_{j\to\infty}v_{j}(x)=\lim_{j\to\infty}\int_{0}^{1}\sigma(\widehat{\gamma}_{xj}(s),Ku_{j-1}(\widehat{\gamma}_{xj}(s),v_{j-1}(\widehat{\gamma}_{xj}(s))),\widehat{\gamma}_{xj}^{\prime}(s))\,{\rm d}s$
$\geq\liminf_{j\to\infty}\int_{0}^{1}\sigma(\gamma(s),Ku_{j-1}(\gamma(s),v_{j-1}(\gamma(s))),\gamma^{\prime}(s))\,{\rm d}s$
again by [20, Thm. 3.23, p. 96], as $\sigma$ is nonnegative, continuous, and convex in $\gamma^{\prime}$ and (4.29) holds.
The curve $\gamma\in{\rm Lip}_{x}$ is hence optimal for $x$, proving the representation formula (2.13).
Step 6: Boundary behavior. Once the representation formula (2.13) is established, we can follow the argument of Step 6 in the proof of Proposition 3.3, by additionally taking into account the dependence of $H$ on $Ku(\cdot,v(\cdot))$. In the frame of the contradiction argument, using (2.14) instead of (3.15) we prove that there exists $\lambda\in(0,1)$ such that
$\gamma^{\prime}(s)\cdot\nabla\varphi(\gamma(s))\leq\lambda\sigma(\gamma(s),Ku(\gamma(s),v(\gamma(s)),\gamma^{\prime}(s))\quad\text{for a.e.}\ s\in(t,1).$ (4.31)
where $\varphi\in C^{1}({\mathbb{R}}^{n})$ is such that $\varphi-v$ has a local maximum at $x\in\partial\Omega\setminus\overline{V_{0}}$, $\gamma\in{\rm Lip}_{x}$ realizes the minimum in (2.13), $t\in(0,1)$, and $\gamma(s)\in B_{r}(x)$ for $s\in[t,1]$ and $r>0$ is fixed but arbitrary. A straightforward adaptation of the chain of inequalities (3.28) yields that $\varphi(\gamma(t))-v(\gamma(t))>\varphi(x)-v(x)$, contradicting the fact that $x$ is a local maximum of $\varphi-v$.
This concludes the proof. ∎
Acknowledgments
This research was funded in whole or in part by the Austrian Science Fund (FWF) projects 10.55776/F65, 10.55776/I5149. Part of this work has been developed during the Thematic Program on Free Boundary Problems at the Erwin Schrödinger International Institute for Mathematics and Physics in Vienna. The warm hospitality of the Institute is gratefully acknowledged. For open-access purposes, the author has applied a CC BY public copyright license to any author-accepted manuscript version arising from this submission.